### v1

# Segment Any Text: A Universal Approach for Robust, Efficient and Adaptable Sentence Segmentation

Abstract
Segmenting text into sentences plays an early and crucial role in many NLP systems. This is commonly achieved by using rule-based or statistical methods relying on lexical features such as punctuation. Although some recent works no longer exclusively rely on punctuation, we find that no prior method achieves all of (i) robustness to missing punctuation, (ii) effective adaptability to new domains, and (iii) high efficiency. We introduce a new model — Segment any Text (SaT) — to solve this problem. To enhance robustness, we propose a new pretraining scheme that ensures less reliance on punctuation. To address adaptability, we introduce an extra stage of parameter-efficient fine-tuning, establishing state-of-the-art performance in distinct domains such as verses from lyrics and legal documents. Along the way, we introduce architectural modifications that result in a threefold gain in speed over the previous state of the art and solve spurious reliance on context far in the future. Finally, we introduce a variant of our model with fine-tuning on a diverse, multilingual mixture of sentence-segmented data, acting as a drop-in replacement and enhancement for existing segmentation tools. Overall, our contributions provide a universal approach for segmenting any text. Our method outperforms all baselines — including strong LLMs — across 8 corpora spanning diverse domains and languages, especially in practically relevant situations where text is poorly formatted. [^1] Our models and code, including documentation, are available at https://huggingface.co/segment-any-text under the MIT license.
[…]
Segment Any Text: A Universal Approach for Robust, Efficient and Adaptable Sentence Segmentation
Markus Frohmann1,2         Igor Sterner3          Ivan Vulić∗3     Benjamin Minixhofer∗3     Markus Schedlthanks: Equal senior authorship.1,2 1Johannes Kepler University Linz   2Linz Institute of Technology, AI Lab 3University of Cambridge markus.{frohmann, schedl}@jku.at,  {is473, bm644, iv250}@cam.ac.uk
2 Introduction
Figure 1: F1 scores and inference time for the prior SoTA (WtP) and our models (SaT and SaT+SM), evaluated on the Ersatz sentence segmentation benchmark. We average over all 23 languages and show the average time (10 runs) for variants with different sizes (L = #layers) to segment 1,000 sentences using consumer hardware (1 Nvidia GTX 2080 Ti GPU).
[…]
Figure 2: Examples of our model’s predictions from (i) ASR output, (ii) multilingual text, and (iii) verse segmentation. (i) shows part of a transcribed TED talk, demonstrating our method is agnostic to punctuation and casing. (ii) is from a Reddit post of German-English translations; existing rule-based systems would segment at nearly every punctuation, and existing neural systems are too reliant on punctuation or need a language code. (iii) shows segmentation of lyrics into verses, showing our model’s predictions in a distinct domain.
Sentence segmentation is defined as the task of identifying boundaries between sentences in a given text. High-quality sentence boundaries are crucial in many NLP tasks and systems since models often expect individual sentences as input (Reimers and Gurevych, 2019, 2020; Liu et al., 2021; Tiedemann and Thottingal, 2020, inter alia). Further, errors in segmentation can have detrimental effects on downstream task performance, e.g., in machine translation (Minixhofer et al., 2023; Wicks and Post, 2022; Savelka et al., 2017).
Existing sentence segmentation tools predominantly rely on punctuation marks. This limitation renders them impractical for text lacking punctuation. To address this issue, some recent methods aim to overcome this dependency (Honnibal et al., 2020; Minixhofer et al., 2023). Specifically, during the training of their model, WtP (Minixhofer et al., 2023) randomly removes punctuation characters to increase robustness against missing punctuation.
However, the performance of WtP as the current state-of-the-art (SoTA) model and all other segmenters is still poor on texts from more challenging domains. This includes, among others, user-generated text such as tweets and highly heterogeneous domains such as lyrics. Segmenting these texts is challenging because of missing and/or extra punctuation, inconsistent spacing, and especially irregular casing. Furthermore, nearly all existing systems, including WtP, require the specification of the texts’ language at inference time. This necessitates an additional preprocessing step of language identification, which often proves to be imperfect, particularly with user-generated content (Lui and Baldwin, 2014; Sterner and Teufel, 2023). Moreover, this necessity limits their applicability to code-switched text, as assigning a single language code to mixed-language inputs is impossible.
To address these challenges, we present a sentence segmentation method that does not rely on language codes or punctuation marks, making it universally applicable across a broad range of languages, corpora, and domains. Specifically, we train subword-based multilingual encoder language models (LMs) in a self-supervised way to predict naturally occurring newlines on web-scale text. We then continue training models on sentence-segmented data in a second, supervised stage to further improve sentence segmentation performance.
We deal with several major issues with previous tools: To ensure robustness against missing punctuation and noise, we propose a set of corruptions, applied randomly to the input during training. Crucially, our method does not rely on language codes. In addition, we mitigate issues observed with short sequences via a novel limited lookahead mechanism. Furthermore, we recognize the variability of sentence boundaries across domains and sentence definitions. To address this, we show how our models can be efficiently adapted to target domains via LoRA (Hu et al., 2022), outperforming previous adaptation methods, especially in data-constrained settings. Further, we improve efficiency by shedding the upper layers of the base model for our default 3-layer models, which segments 1000 sentences in approx. $0.5$ seconds on our hardware.
Figure 1 shows that the standard 3-layer version of SaT outperforms the current state-of-the-art, WtP, while achieving a $\approx 3$x reduction in inference time. Overall, we present several innovations that overcome each of the shortcomings of previous methods, culminating in a universal model for sentence segmentation. We provide some examples of our model’s predictions in Figure 2.
Contributions.
1) We introduce Segment any Text (SaT), an efficient method for sentence segmentation that can reliably segment text across 85 languages regardless of lexical features such as punctuation or casing. 2) We show how our models can be adapted to different domains via data-efficient means, requiring only a minimal set (e.g., 16) of sentence-segmented examples. 3) We train and release SaT models in five sizes, covering 85 languages, and demonstrate state-of-the-art performance across 8 corpora, even outperforming newly introduced strong LLM baselines.
3 Background and Related Work
We start by providing an overview of existing sentence segmentation systems. Following Read et al. (2012), we categorize them into 1) rule-based, 2) supervised statistical, and 3) unsupervised statistical approaches. Then, we discuss the recently introduced state-of-the-art approach, WtP. Moreover, we discuss domain-specific segmentation approaches. Lastly, since we are the first to evaluate large language models (LLMs) for sentence segmentation broadly, we briefly survey them and discuss their usage in sentence segmentation tasks.
3.1 General Systems and Baselines
1. Rule-based methods segment text into sentences using hand-crafted rules. The segmenters in Moses (Koehn et al., 2007) and SpaCy (Honnibal et al., 2020) split on punctuation characters, except for predefined exceptions like abbreviations and acronyms. PySBD (Sadvilkar and Neumann, 2020) relies on exceptions and regular expression rules. Although generally efficient, these methods demand manual per-language effort to incorporate language-specific rules. This also necessitates specifying a language code at inference time.
2. Supervised statistical methods learn segmentation from a sentence-segmentation annotated corpus. One early method by Riley (1989) involved a decision tree to determine if each punctuation mark in a text represents a sentence boundary based on linguistic features surrounding punctuation. Satz (Palmer and Hearst, 1997) and Splitta (Gillick, 2009) build on this approach but utilize neural networks and SVMs, respectively. Similarly, in Ersatz, Wicks and Post (2021) propose to use a Transformer (Vaswani et al., 2017) with subwords as context around punctuation marks. However, these methods are limited by their reliance on punctuation to define sentence boundaries. This becomes problematic in poorly punctuated texts as non-punctuation characters cannot serve as sentence boundaries. Breaking from this limitation, the dependency parser in the SpaCy library (Honnibal et al., 2020) jointly learns dependency parsing and sentence segmentation on a labeled corpus without special treatment of punctuation.
3. Unsupervised statistical methods predict sentence boundaries from unsegmented text alone. Kiss and Strunk (NLTK; 2006) use features such as character length and internal punctuation to identify abbreviations, initials, and ordinal numbers, treating all other punctuation as sentence boundaries. Furthermore, Wicks and Post (2021) additionally introduces an unsupervised version of Ersatz, relying on punctuation preceding paragraph breaks.
3.2 Where’s the Point (WtP)
WtP represents the current state-of-the-art in sentence segmentation (Minixhofer et al., 2023). Like our method, it can be used in unsupervised and supervised variations. Hence, we choose WtP as our main baseline and examine it in the following.
[…]
Robustness to corruptions. To make WtP less reliant on punctuation, Minixhofer et al. (2023) randomly remove some punctuation during training. In addition, they predict the likelihood of commonly occurring punctuation as an auxiliary objective. While this helps make WtP models less reliant on punctuation, we still find that WtP models have major issues when text is inconsistently formatted, especially irregular casing.
Efficiency. WtP uses the character-level encoder LM Canine-S (Clark et al., 2022) as its backbone. Operating on characters as the fundamental unit constitutes a major bottleneck in terms of speed, resulting in poor efficiency.
Multilinguality. To increase language-specific capacity, WtP utilizes language adapters (Pfeiffer et al., 2022). This, however, confines its multilingual abilities since a language code must be specified at inference time. This is especially problematic in code-switching, where multiple languages are present, leading to ambiguity.
Short texts. We also found WtP models deficient in segmenting short sequences, such as tweets or sentence pairs. During training, paragraphs are packed to always fully use the model’s context size. While being efficient at training, we hypothesize that this renders short sequences out-of-domain.
Domain adaptation. Minixhofer et al. (2023) also evaluate two supervised adaptation methods. First, WtPT tunes the segmentation threshold $\alpha$ based on an already sentence-segmented corpus. Second, based on the auxiliary punctuation prediction objective, WtPPUNCT fits a logistic regression on the probability distribution of the punctuation logits. However, these kinds of adaptations fall short on more challenging domains such as lyrics and code-switched text, as quantified later in Section 6.3.
3.3 Domain-specific Sentence Segmentation
Due to deviations from typical sentence structures, differences in sentence lengths, and non-standard punctuation, sentence segmentation is highly dependent on domain-specific characteristics (Sheik et al., 2024), providing a strong basis for domain-specific systems (Read et al., 2012).
Prior studies have focused on creating a dedicated model for a single domain. Reynar and Ratnaparkhi (1997) utilized features unique to the financial domain. Tuggener and Aghaebrahimian (2021) hosted a shared task on transcripts of spoken texts. Brugger et al. (2023) train models to segment sentences in the legal domain.
Previous approaches to segmenting lyrics into verses require songs to be already pre-segmented into lines. Watanabe et al. (2016) extract features based on repeated patterns and part-of-speech. Fell et al. (2018) improve upon this approach by using convolutions and a more refined set of features.
In contrast to prior domain-specific models, we propose a single model that can be efficiently adapted for segmenting sentences from wildly heterogeneous domains and languages, outperforming previous domain-specific models, even when using only a limited number of examples.
3.4 Large Language Models
Large language models (LLMs) have become a de facto tool for use in many NLP tasks (Zhao et al., 2023; Minaee et al., 2024). Most modern LLMs are decoder-only Transformers (Vaswani et al., 2017; Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2024, inter alia). Recently, prompting has emerged as the dominating paradigm for solving a task (Ouyang et al., 2022; Liu et al., 2023).
However, despite widespread use, LLMs have yet to be extensively evaluated for sentence segmentation. In this work, we aim to bridge this gap by shedding light on how well popular LLMs can segment sentences when prompted to do so, particularly in more challenging domains such as lyrics, where using LLMs may be especially valuable.
4 SaT: Segment any Text
To create a reliable and effective system across various scenarios, we pre-train a model on paragraph segmentation as in Minixhofer et al. (2023). In the following, we outline how we solve each of the major issues of WtP discussed earlier, leading to a universal model for sentence segmentation.
Efficiency. We resort to models using subword tokenization, processing tokens consisting of multiple characters at a time, making them considerably faster than their character-level counterparts.
Multilinguality. Unlike Minixhofer et al. (2023), we do not rely on language adapters. In addition to improving inference time and storage requirements, this also improves multilinguality since no language has to be specified at inference time.
Robustness to corruptions. We randomly remove common punctuation-only tokens with probability $p$ and use the auxiliary punctuation-prediction objective during training. Further, we randomly remove all casing and punctuation in 10% of samples within a batch during training. The resulting model, Segment any Text (SaT), already shows strong segmentation performance at improved efficiency.
Still, to further improve SaT, we continue training it on a Supervised Mixture of already-segmented sentences. [^2] Although more domain-specific, this is motivated by the strong multi-task learning abilities of T5 (Raffel et al., 2019). To be even less dependent on patterns such as punctuation, spaces, and casing, we augment the data by introducing several additional corruption schemes, resulting in our more specialized model, SaT+SM.
Our first corruption scheme removes all casing and punctuation tokens for all text. Secondly, we add randomness to corruption in as many situations as we deem useful, e.g., user-generated text in tweets or forums. This includes duplicating punctuation, removing punctuation, lowercasing, and removing/adding spaces between sentences. Finally, we also use clean, non-corrupted text. We then sample uniformly across these three categories. For details, see Appendix A.2.
Domain Dataset Description Characteristics Source
Clean Text Universal Dependencies (UD) Treebanks in many languages. Includes gold-standard segmentation into sentences. de Marneffe et al. (2021); Nivre et al. (2020)
OPUS100 Sentences from subtitles and news in 100 languages. A challenging sentence segmentation benchmark (Zhang et al., 2020) Tiedemann (2012)
Ersatz Sentences from WMT Shared Tasks, mainly comprising news (commentary). Includes manual sentence segmentation corrections by Wicks and Post (2021). Wicks and Post (2021); Barrault et al. (2020)
Noisy Text SEPP-NLG Shared Task (surprise test set) 500 transcribed public TED talks in each of 4 European languages. Neither casing nor punctuation tokens are present. Tuggener and Aghaebrahimian (2021)
Tweets User-generated content in the form of Slovene (sl) and Serbian (sr) tweets. Noisy; short in length (70/115 characters on average for sl and sr, respectively). Fišer et al. (2020); Miličević and Ljubešić (2016)
Code- switching C.f. Table 10. Reddit posts for German-English (de-en); data for 3 additional language pairs is taken by concatenating code-switching sentences from bilingual transcriptions. We treat all data as transcriptions, removing all punctuation and casing; we only keep sentences with at least one token of each language. Deuchar (2009), Osmelak and Wintner (2023), Nguyen and Bryant (2020) and Çetinoğlu (2017)
Legal MultiLegalSBD Laws and judgements from legal documents in 6 languages. Domain-specific jargon and structure; formal and complex sentences. Brugger et al. (2023)
Lyrics Verses 35,389 English songs across 16 genres spanning 3 levels of repetitiveness. We replicated the setup by Fell et al. (2018). Meseguer-Brocal et al. (2017)
Table 1: Overview of the evaluation corpora we use. For more details, see Appendix A.2.
Short texts. To resolve issues with short sequences, we enforce SaT to use only the immediate $N$ future tokens for its predictions. We do so via a limited lookahead mechanism. Let $k_{i}$ be the token occurring at position $i$, and $\mathbf{a_{ij}}$ its corresponding attention mask, where $j$ corresponds to the token to be attended to. A naive modification of the attention mask would set $\mathbf{a_{ij}}=0$ for $j>i+N$. However, using Transformer networks with multiple layers results in a lookahead of $N\times L$, where $L$ is the number of Transformer layers (Jiang et al., 2023). We thus split up the lookahead evenly into $L$ layers, resulting in the following attention mask:
$\mathbf{a}_{\mathbf{ij}}=0\text{ for }j>i+N_{L},$
where $N_{L}$ is the per-layer lookahead, i.e., $N_{L}=\frac{N}{L}$. Using an intermediate value for $N$ makes SaT robust to both short and long sequences – relying on some future context where appropriate, but not so much that it falters on short sequences.
Limited lookahead can be thought of as sliding window attention (Beltagy et al., 2020; Jiang et al., 2023) with two crucial tweaks: 1) the sliding window extends forward into the future instead of backward, 2) past tokens are not masked out.
Domain adaptation. Finally, some domains require more sophisticated adaptation than only changing the threshold or relying on punctuation logits. We thus explore low-rank adaptation (LoRA; Hu et al., 2022) to adapt our models efficiently, denoted by SaT+LoRA. We show how this enables state-of-the-art performance on verse segmentation using our models later in § 6. In our setup, it trains $\approx 1\%$ of the parameters of SaT but results in no inference overhead since LoRA weights can be merged into the backbone LM weights at inference time (Pfeiffer et al., 2023).
5 Experimental Setup
5.1 Evaluation
To evaluate our method, we compare ground truth and predicted sentence boundaries on the test sets of corpora spanning a diverse set of languages, sources, and domains, [^3] We acknowledge the concept of domains remains an open issue in NLP (Holtermann et al., 2024; Raffel et al., 2019). summarized in Table 1.
In addition, to evaluate how well our method can segment short sequences, we generate non-overlapping sentence pairs from the datasets categorized as clean text. We additionally simulate a real-time automatic speech recognition (ASR) scenario using transcripts from speeches in 76 languages. We generate sentence pairs in a similar way and remove all punctuation as well as all casing.
We report character-level F1 scores for the positive (i.e., sentence-ending) labels. For short sequences, we use the proportion of perfectly segmented sequences within a corpus; this is stricter than F1 since any segmentation error results in a score of zero for the entire sequence. For SEPP-NLG, we use the evaluation script and surprise test set provided by the shared task organizers (Tuggener and Aghaebrahimian, 2021), reporting F1 scores on the token level. In our evaluations on clean text across all 85 languages, we run all competitor and baseline systems ourselves. For these results, we test all differences for significance with paired two-tailed permutation tests. We approximate them with $N\!=\!10,000$ and set the significance threshold at $\alpha\!=\!0.05$. Additional evaluation and dataset details are provided in Appendix A.2.
Model ar cs de en es fi hi ja ka lv pl th xh zh 81 langs
Multilingual
SpaCyM - 91.1 84.7 91.5 94.5 93.5 - - - 91.4 94.0 - - - -
Ersatz 77.2 90.9 87.0 91.4 95.1 93.9 84.8 69.3 - 91.1 94.8 - - 77.2 -
Llama 38B 78.2 93.4 92.6 95.2 96.0 95.5 85.6 64.7 89.2 93.0 96.2 66.0 71.7 82.0 79.1
Command R 58.6 68.1 79.1 84.6 81.0 74.1 72.0 52.2 25.6 74.2 78.6 10.6 56.1 73.7 55.6
SaT 79.9 91.7 90.4 93.6 94.0 94.2 84.9 88.6 75.7 92.2 93.7 68.0 80.3 78.0 84.9
SaT+SM 80.7 95.7 94.0 96.5 97.3 96.9 90.3 88.1 93.6 96.1 97.7 72.9 89.6 88.9 91.6
Monolingual
NLTK - 90.8 87.1 92.2 94.1 93.9 - - - - 94.5 - - - -
PySBD 37.4 - 80.6 69.6 56.9 - 70.1 76.1 - - 49.3 - - 86.9 -
SpaCyDP - - 89.0 92.9 93.5 94.1 - 77.1 - - 95.3 - - 87.7 -
WtP 77.3 91.1 89.2 93.9 93.2 93.4 85.0 72.7 91.3 90.4 93.6 66.6 77.2 90.7 84.2
WtPT 79.9 92.0 92.0 93.5 94.2 94.1 85.2 85.6 91.1 93.1 93.5 69.7 80.7 89.3 85.9
WtPPUNCT 85.4 96.4 95.0 96.7 97.4 97.7 90.8 93.1 92.8 96.6 97.5 71.3 89.8 95.5 91.7
SaT+LoRA 86.3 96.2 95.4 96.7 97.7 97.5 92.9 94.4 93.3 97.0 97.7 73.7 90.8 94.9 93.1
Table 2: Mean sentence segmentation F1 scores over OPUS100, UD and Ersatz. For the average, we report macro F1 over languages from all datasets where train and test sets are available. Results are shown using 3-layer variations of all models. Numerically best results are in bold, statistically indistinguishable ones from this best are underlined.
Model en de fr it Avg.
htw+t2k 77 82 76 75 78
OnPoint 80 82 77 77 79
Unbabel 83 78 78 76 79
SaT 73.4 79.9 73.1 72.9 74.8
SaT+SM 79.7 84.0 78.3 77.1 79.8
Table 3: F1 scores on the surprise test set of the SEPP-NLG Shared Task. For comparison, we provide results for the 3 best-performing systems from the Shared Task. We use 12-layer versions of our models.
Baselines. We compare against PySBD and NLTK as representatives of rule-based and unsupervised statistical methods. For supervised methods, we evaluate the punctuation-agnostic SpaCyDP and Spacy’s multi-language model, SpaCyM. We also compare against Ersatz. Our main comparison is against the current SoTA models: WtP, WtPT, and WtPPUNCT.
LLM-based baselines. To evaluate LLMs, we use 1) Cohere’s Command R as a recent LLM with claimed strong multilingual performance, and 2) Meta’s Llama 38B due to its popularity and strong performance. Officially, Command R supports 23 languages, whereas Llama 38B only supports English. [^4] Due to imperfect filtering of common web-crawled corpora, all LLMs can be considered multilingual to some extent. We split up each dataset into chunks of $10$ sentences to avoid cases where sentences are cut off at critical positions and observed issues with long context lengths. Then, we prepend the prompt to each chunk and let the LLM segment $10$ sentences. [^5] For a fair comparison, we thus exclude every 10th label when calculating F1 scores. For songs and short sequences, we feed in whole samples and hence do not exclude any labels. Finally, to make evaluation metrics robust to unwanted alterations of the input by the LLM, we apply the Needleman-Wunsch algorithm (NW; Needleman and Wunsch, 1970) to align sentences within each input and output chunk. For the prompt and other implementation details, including alignment via NW, we refer to Appendix A.2.
5.2 Training Setup
We train Transformer models operating on subwords, initialized with the weights of XLM-RoBERTa (XLM-R; Conneau et al., 2020). We use a lookahead limit of $48$ tokens, which we found to work well in practice on text of any length, leading to SaT. We use the mC4 (Raffel et al., 2019) corpus and sample text uniformly from the 85 languages also used by Minixhofer et al. (2023).
To train SaT+SM, we continue training SaT on the training set of UD due to its high quality and availability in most of the 85 considered languages. For languages without UD data, we resort to silver-quality data from OPUS100 or NLLB (Costa-jussà et al., 2022), whichever is available.
To adapt to different user requirements w.r.t. efficiency, we train and release SaT and SaT+SM models in different sizes from 1-12 layers, where we remove the upper layers for models $<12$ layers.
For adaptation via LoRA (SaT+LoRA), we use SaT as a starting point. [^6] We include its task head since we found that it improves stability. We also experimented with applying LoRA to SaT+SM, but did not find it to improve upon SaT+LoRA. We use the respective training set using max. $10,\!000$ sentences.
The full details of the experiment setup regarding the datasets, infrastructure, training, and hyperparameters are provided in Appendix A.2.
6 Results
6.1 Performance on Clean Text
Table 2 shows evaluation results on clean text, averaged over OPUS100, UD, and Ersatz on a diverse selection of languages, including an average over 81 languages.777 For the average, we only consider languages with datasets with both train and test sets for a fair comparison. While we evaluate on 85 languages, this is the case in 81 languages. We categorize methods into (i) multilingual, which take only text as input, and (ii) monolingual, which additionally rely on a language code or, in the case of WtPT, WtPPUNCT, and SaT+LoRA, are adapted to a target domain.
Both SaT and SaT+SM outperform the current non-domain-adapted SoTA model, WtP. Meanwhile, unlike WtP, our models do not rely on specifying a language code as input.
Remarkably, SaT+SM and WtPPUNCT are not statistically significantly different, achieving average F1 scores of $91.6$ and $91.7$ respectively. This is despite WtPPUNCT relying on adaptation to a target sentence-segmented corpus, whilst SaT+SM is a general-purpose multilingual model. Finally, SaT+LoRA significantly outperforms the existing domain-adapted SoTA, WtPPUNCT, making it the best overall model. Our domain-adapted model outperforms WtPPUNCT in 63 out of 81 languages.
Among the LLMs, Command R, despite being trained in 23 languages, does surprisingly poorly, with Llama 38B surpassing it by $23.5\%$ absolute avg. F1. Nevertheless, Llama 38B still falls short compared to all variants of SaT. On the English benchmarks, given the abundance of English text, we expected our models to be easily outperformed by LLMs; yet, unlike WtP, SaT+SM outperforms both LLMs on every dataset.
We provide full per-dataset results, including all 85 languages, in § A.4. We also conduct ablation studies on each of our stages’ components in § A.1.
Model Tweets Sentence Pairs Macro Avg.
sl sr Speeches Ersatz
Llama 38B 73.4 76.0 66.9 94.8 77.8
Command R 53.8 47.4 23.0 70.0 48.6
WtP 70.8 71.4 12.6 78.0 58.2
WtPT 70.4 71.4 18.9 79.0 59.9
WtPPUNCT 80.1 82.3 37.9 91.5 72.9
SaT 80.5 75.5 28.8 84.0 67.2
SaT+SM 78.0 72.9 41.7 92.5 71.3
SaT+LoRA 87.2 89.1 56.8 93.9 81.8
Table 4: Proportion of perfectly segmented short sequences. For Speeches and Ersatz, we are averaging scores over languages. We use 12-layer versions of each model given the task’s increased difficulty. [^9] We exclude other baselines since none of them support sl/sr or all languages from TED/Ersatz.
6.2 Performance on Noisy and Short Text
Table 3 presents the results of our method when evaluated on the SEPP-NLG Shared Task. SaT+SM establishes a new state-of-the-art, outperforming the SEPP-NLG winners. This is despite our model supporting 81 additional languages and use cases not considered in the Shared Task.
Furthermore, Table 9 shows evaluation results on short sequences, including tweets and sentence pairs taken from manually corrupted speeches and Ersatz. We observe similar patterns on these corpora: SaT and SaT+SM outperform WtP, improving avg. F1 scores by $9\%$ and $13.1\%$, respectively, SaT+LoRA continues to be the best overall model, also outperforming both LLMs. We additionally provide an ablation study showing the importance of limited lookahead in SaT in Table 9.
Model es en de en vi en tr de Macro Avg.
Llama 38B 47.9 56.3 35.5 33.9 43.4
Command R 30.4 51.9 30.0 17.6 32.5
SpaCyDP * 17.6 8.6 11.3 12.2 12.2
WtP * 38.6 39.0 25.5 33.5 29.1
WtPT * 52.2 45.7 46.7 34.4 43.2
WtPPUNCT * 62.1 60.1 59.0 41.0 54.9
SaT 54.5 49.2 49.3 39.8 48.2
SaT+SM 59.6 58.4 57.3 42.4 54.4
SaT+LoRA 65.0 65.6 67.5 48.8 61.7
Table 5: Sentence segmentation F1 scores for code-switched text. We use 12-layer versions of each model. * indicates models using language codes, where we try both language codes and show the better score. We show results using both language codes in Appendix A.4.
6.3 Performance on Challenging Domains
Code-switching. The results in Table 5 reveal that WtP achieves an average F1 score of $29.1\%$, while the highest-performing LLM scores $43.4\%$. SaT and SaT+SM achieve average F1 of $48.2\%$ and $54.4\%$, respectively. SaT+LoRA continues to improve performance, achieving $61.7\%$. To the best of our knowledge, this is the first evaluation of sentence segmentation on code-switching text. While our models represent SoTA, the evaluation results indicate that it is a challenging task.
We now evaluate domain adaptation performance of our method on two highly distinct domains: lyrics and legal data.
Lyrics. Table 6 shows results on verse segmentation (i.e., segmenting songs into verse, chorus, etc.). None of the other baseline systems, including LLMs, can improve over the current domain-specific SotA, $\text{SSM}_{\text{string}}$. In contrast, SaT+LoRA outperforms $\text{SSM}_{\text{string}}$ by $10\%$ avg. F1. The difference is especially pronounced in hard-to-segment songs that are low in repetitiveness (e.g., Rap music), with a $15\%$ difference in F1 scores. When evaluating SaT+LoRA on manually corrupted lyrics, it still outperforms all baselines, even when compared to baselines evaluated on non-corrupted songs. Additionally, SaT+LoRA @1000, using 1000 songs per genre for adaptation, still outperforms all baselines. We provide complete results, including those for each genre, in Appendix A.4.
Legal. We provide comprehensive results on MultiLegalSBD in Appendix A.4.
Model Corrupted? Repetitiveness
High Mid Low
$\text{SSM}_{\text{string}}$ 63.8 71.3 64.8 47.3
Llama 38B 45.5 49.7 48.9 46.7 33.8
Command R 36.3 38.3 38.0 37.1 28.7
WtPPUNCT @100 46.9 53.8 55.8 55.2 35.9
WtPPUNCT @1000 49.1 56.1 58.4 57.5 44.9
WtPPUNCT 49.2 56.2 58.4 57.6 44.9
SaT+LoRA @100 60.8 62.4 67.8 62.9 51.6
SaT+LoRA @1000 67.3 72.4 76.5 73.1 62.7
SaT+LoRA 68.5 73.8 77.9 74.8 62.3
Table 6: Macro-averaged verse segmentation performance over per-genre F1 scores. Values for $\text{SSM}_{\text{string}}$ taken from Fell et al. (2018), with lyrics already pre-segmented into lines. @N corresponds to using a maximum of N songs per genre for adaptation.
7 Discussion
LLMs. Contrary to our expectations, our evaluation results reveal that LLMs generally underperform, particularly in non-English languages. Notably, when using LLMs for sentence segmentation via prompting, each sentence is processed twice – once as part of the input, appended to the prompt, and once within the output. This redundancy leads to inefficient processing, needing to copy the input verbatim to the output, ideally only adding newlines. However, in practice, LLMs are highly prone to alter their input (Barbero et al., 2024). We found this issue to be particularly severe for noisy text and lyrics. [^10] Llama 38B and Command R altered $1.5\%$ and $2\%$ of all characters within lyrics, respectively, even though we prompted them not to alter their input (c.f. Appendix A.2). This is highly problematic for a specific task requiring input and output characters to remain the same. Still, we tried to address this by using the Needleman-Wunsch sequence alignment algorithm to make pure segmentation performance comparable to other methods.
Aiming to improve the segmentation performance of LLMs, we experimented with few-shot prompting. However, this did not yield the desired improvements; in fact, it degraded performance. Additionally, we tested varying the number of input-output sentences. The results of both of these ablation studies are presented in Appendix A.1.
Figure 3: Macro avg. F1 vs. number of sentences used for adaptation, averaged over languages in {OPUS100, UD, Ersatz}. Per-corpus results shown in Appendix A.1.
Efficiency. For our method, we rely on XLM-R as the LM backbone. Operating on subwords makes SaT considerably faster than WtP. We compare sentence segmentation performance and time on Ersatz across model sizes from 1-12 layers, illustrated in Figure 1. Additional datasets are shown in Figure 4. The standard 3-layer variations of SaT take $\approx 0.5$ seconds to segment 1000 sentences on the hardware specified in Appendix A.2, making them $3$ times faster than WtP, while also outperforming WtP models in all sizes on Ersatz. Furthermore, for SaT, performance plateaus with sizes $>3$ layers, whereas SaT+SM continues to improve when scaling up its size, making it by far the best model, despite never being exposed to Ersatz.
Few-shot domain adaptation. We now analyze how many sentences are needed to adapt our domain adaptation method, SaT+LoRA, to a target corpus, and compare it to previous methods. As shown in Figure 3, WtPT and WtPPUNCT perform similarly when using $1024$ sentences for domain adaptation. However, WtPPUNCT fails to outperform the fully self-supervised variation of WtP when $\leq 32$ sentences are available. In contrast, SaT+LoRA markedly improves upon the self-supervised SaT with only $16$ available sentences, and outperforms WtPPUNCT by almost $10\%$ in F1 score, making it substantially more sample-efficient.
8 Conclusion
We proposed SaT, an efficient, robust, and highly adaptable multilingual sentence segmentation method that neither relies on language codes nor punctuation. Further, we introduced SaT+SM, improving SaT via supervised adaptation using multiple corruption schemes. Our method consistently achieves state-of-the-art performance in experiments across 85 languages and 8 diverse corpora, even outperforming newly introduced and optimized strong LLM baselines. We also demonstrated that SaT can be efficiently domain-adapted via LoRA, setting new performance standards on segmentation of lyrics and code-switching text. Overall, we hope SaT will unlock significantly improved text data (pre-)processing across a range of NLP applications for multiple languages and domains via its robust and consistently strong performance, versatility, and high efficiency.
Limitations
To the best of our knowledge, our evaluations are the most comprehensive to date, spanning 8 diverse corpora across different domains, languages, and noise levels, and sequence lengths. Still, we may not have covered every possible scenario. Second, since we use XLM-R as our backbone, we also use its tokenizer, which has been shown to tokenize text less efficiently in some language (Liang et al., 2023), potentially exacerbating existing biases. We try to minimize bias w.r.t. performance by sampling text from all languages uniformly in both stages. Furthermore, our use of subword LMs merges characters into subwords. Theoretically, this could limit sentence boundaries to end-of-token positions; however, in practice, we did not find this to be an issue. Finally, language support could be further improved by e.g., replacing mC4 with MADLAD-400 (Kudugunta et al., 2023) for the pre-training stage. We leave this to future work.
Ethical Considerations
Our work is multifaceted, as are the ethical dimensions it encompasses. First, we acknowledge the possibility of NLP datasets and models for encoding unfair stereotypical (Blodgett et al., 2020) and exclusive (Dev et al., 2021) biases that may lead to representational and allocational harms (Barocas et al., 2017). This issue is a general property of pre-trained LMs, and the models and datasets utilized in our study are similarly at risk. We advise practitioners to use these models with the appropriate care and we refer to existing works (Liang et al., 2021; Lauscher et al., 2021) for discussions on bias mitigation. Second, one key aspect of our work deals with efficiency. On the one hand, considering the well-documented relationship between model training efforts and potential CO2 emissions (Strubell et al., 2019), our research contributes to Green AI by improving the environmental sustainability of state-of-the-art sentence segmentation systems. On the other hand, since the training of language models often comes with high infrastructure prerequisites only available to certain user groups (Bender et al., 2021), we hope that our work also contributes to the continued democratization of language technology by reducing resource- and language-related usage barriers.
Acknowledgments
This research was funded in whole or in part by the Austrian Science Fund (FWF): P36413, P33526, and DFH-23, and by the State of Upper Austria and the Federal Ministry of Education, Science, and Research, through grants LIT-2021-YOU-215. In addition, Ivan Vulić and Benjamin Minixhofer have been supported through the Royal Society University Research Fellowship ‘Inclusive and Sustainable Language Technology for a Truly Multilingual World’ (no 221137) awarded to Ivan Vulić. This research has also been supported with Cloud TPUs from Google’s TPU Research Cloud (TRC). This work was also supported by compute credits from a Cohere For AI Research Grant, these grants are designed to support academic partners conducting research with the goal of releasing scientific artifacts and data for good projects. We also thank Simone Teufel for fruitful discussions.
Appendix A Appendix
(a) OPUS100
(b) UD
(c) Code-Switching
Figure 4: F1 scores and inference time for the prior SoTA (WtP) and our models (SaT and SaT+SM), evaluated on additional sentence segmentation benchmarks. We average over all 23 languages and show the average time (10 runs) for variants with different sizes (L = #layers) to segment 1,000 sentences using consumer hardware (1 Nvidia GTX 2080 Ti GPU). Performance on Ersatz is shown in Figure 1.
A.1 Ablation Studies
(a) Effect of few-shot prompting.
(b) Effect of varying the number of input-output sentences.
Figure 5: Ablation study on sentence segmentation performance of LLMs.
Model Variation Clean Text Tweets Code Switching
SaT - 84.9 78.0 48.2
Only clean text 84.7 33.5 16.5
SaT+SM - 91.6 75.5 54.4
Only clean text 91.5 77.2 10.2
No pre-training 89.9 42.1 44.0
SaT+LoRA - 93.1 88.2 61.7
No pre-training 88.4 74.5 12.4
Table 7: Effect of various components of our method’s variants. We report macro average F1 scores for each domain and use models with the same number of layers for each category of text as in the main text. Only clean text does not apply any corruptions. No pre-training skips the paragraph segmentation stage on web-scale mC4, and thus starts from XLM-R weights. Best per-category results are bold.
Model Variation Lyrics Legal
Corrupted? Corrupted?
SaT+LoRA @100 - 60.8 62.4 81.1 93.6
No pre-training 20.3 34.3 61.2 79.8
SaT+LoRA - 68.5 73.8 83.3 95.1
No pre-training 59.9 62.1 82.5 94.9
Table 8: Effect of the web-scale pre-training stage on adaptation to hard domains, averaged over genres/legal categories. @100 corresponds to using a maximum of 100 songs/documents per genre/category for adaptation.
SaT components.
We show the effect of removing different components of our corruption schemes used in SaT and SaT+SM in Table 7. For SaT, only using clean text even slightly hurts performance on clean text and strongly degrades performance in our more noisy tweets and code-switching evaluations. A similar pattern occurs for SaT+SM: Only using clean text hurts performance. Moreover, skipping the web-scale pre-training stage (No pre-training) also decreases performance to a large extent, with the difference being particularly large for tweets and code-switching. Finally, for SaT+LoRA, no pre-training similarly degrades performance, with the difference being particularly large in code-switching, where SaT+LoRA is better by $49.3\%$ absolute F1.
Moreover, Table 8 compares domain adaptation performances via LoRA to models without our web-scale pre-training to SaT models with it. As observed before, no pre-training markedly degrades performance in both lyrics and legal data. The difference is especially large in cases where only 100 songs or documents are available, clearly showing that our pre-training stage improves sample-efficiency.
Limited lookahead.
We further provide an ablation study on the effect of disabling the limited lookahead mechanism using sentence pairs in Table 9. Without limited lookahead, SaT is outperformed by WtP. On the contrary, with limited lookahead, SaT outperforms WtP by a considerable margin, where the difference is even more pronounced for 12-layer variations. Moreover, SaT+SM hardly benefits from limited lookahead, justifying our decision to disable it for SaT+SM.
Effect of model size.
Figure 4 shows the effect of scaling up model sizes on OPUS100, UD, and code-switching, respectively. Remarkably, all 3-layer variations of SaT+SM clearly outperform WtP, despite not relying on language codes and being $\approx 5$x faster. The difference is particularly pronounced in code-switching, where even the 1-layer variations of both SaT and SaT+SM outperform the best variation of WtP. In general, performance continues to increase when further scaling up model sizes up to 12 layers.
Model Layers Look- ahead OPUS100 UD Ersatz TED
WtP 3 $\infty$ 52.4 80.6 78.0 9.8
12 $\infty$ 52.8 77.9 78.0 12.6
SaT 3 $\infty$ 4.4 1.8 3.5 1.9
48 56.9 82.4 82.2 16.9
12 $\infty$ 31.0 55.0 55.5 20.9
48 63.3 85.2 84.0 28.8
SaT+SM 3 $\infty$ 72.2 91.4 85.9 29.4
48 73.7 93.1 85.8 28.4
12 $\infty$ 78.0 93.5 92.5 41.7
48 78.6 93.6 91.3 38.3
Table 9: Proportion of perfectly segmented sequences within additional corpora.
LLMs.
Figure 5 shows the effect of few-shot prompting and varying the number of input-output sentences for Llama 38B and Command R. Contrary to our expectations, in-context learning via few-shot prompting does not improve sentence segmentation performance for both LLMs in consideration. Providing only a single example already degrades performance, and providing more examples further degrades it. Furthermore, increasing the number of input-output sentences from our favorably low default of $10$ results in considerable performance decreases. Notably, when using $80$ input-output sentences per chunk, both LLMs achieve F1 scores of only $<60\%$.
A.2 Complete Experiment Details
Language Number of Sentences Source
Train Test
sl 2728 2728 Fišer et al. (2020)
sr 1727 192 Miličević and Ljubešić (2016)
es-en 1335 1334 Deuchar (2009)
en-de 678 599 Osmelak and Wintner (2023)
tr-de 578 805 Çetinoğlu (2017)
vi-en 1360 1361 Nguyen and Bryant (2020)
Table 10: Number of train and test sentences from tweets and code-switched text, including their source.
Repetitiveness Genre Number of Songs
Train Test
High Punk Rock 778 190
Pop Punk 512 141
Country 2916 711
Mid Rock 4611 1182
Pop 3490 891
RnB 3542 915
Alternative Rock 3370 856
Alternative Metal 651 155
Soul 494 110
Hard Rock 1821 430
Indie Rock 1193 305
Pop Rock 1633 412
Heavy Metal 988 216
Indie Rock 1193 305
Low Southern Hip Hop 836 208
Gangsta Rap 270 64
Table 11: Number of train and test songs per genre.
Language Number of Documents
Laws Judgements
Train Test Train Test
de 10 3 104 27
en - - 64 16
es 494 183 151 39
fr 1672 459 252 63
it 2206 704 194 49
Table 12: Number of legal train and test documents per category. We discard Portuguese since there is no training data available.
Dataset details.
We give an overview of all used languages and their evaluation dataset sizes for clean text in Table 15. Furthermore, we provide statistics of splits for noisy text and additional domains in Table 10 for tweets and code-switching, Table 11 for lyrics, and Table 12 for legal data.
If a given corpus does not provide train and test splits, we set aside 10,000 sentences for testing and keep the rest for training if more than 10,000 sentences are available. If a corpus is smaller, we set aside 50% for testing and use the rest for adaptation. To train SaT+SM, if neither UD nor OPUS100 train data is available, we resort to NLLB. This is the case in Cebuano (ceb), Javanese (jv), Mongolian (mn), and Yoruba (yo), where we take 10,000 sentences each.
To simulate our real-time automatic speech recognition (ASR) scenario, we take publicly available TED talk transcripts in 76 languages, available at opus.nlpl.eu/TED2020/corpus/version/TED2020. We generate non-overlapping sentence pairs as done in other experiments to evaluate performance on short sequences. Additionally, we remove fully lowercase all pairs and all punctuation tokens. We generally derive punctuation tokens with the commonly used Moses tokenizer (Koehn et al., 2007) for languages where it is available. For all other languages, we simply remove all punctuation characters.
Moreover, we observe that used tweets in sl and sr are inconsistent w.r.t. segmenting single emojis. We thus filter out all emojis as a simple pre-processing step. Similarly, we normalize tweets by filtering out words starting with http, #, and @.
Computing infrastructure.
We train SaT on a TPUv4 VM with 8 cores, SaT+LoRA on a TPUv3 VM with 1 core, and SaT+SM using a single A100 GPU. To measure inference time, we use a consumer-grade GPU, the Nvidia GTX 2080 Ti with an AMD EPYC 7402P CPU.
Implementation details.
We use the PyTorch (Paszke et al., 2019) and transformers (Wolf et al., 2020) libraries for all experiments. For adaptation via LoRA, we make use of the adapters library (Poth et al., 2023; Pfeiffer et al., 2020) library, a wrapper around the transformers library. Our code and models are released under the MIT License, ensuring open access to the community for further development.
Training of SaT.
We train SaT using a context window of $256$ since we observed that it improves performance. During inference, we use the full context size of XLM-R, $512$. Moreover, we follow Minixhofer et al. (2023) and sample paragraphs to ensure that a maximum of 10% of paragraphs do not end in punctuation (except for Thai, which does not use sentence-ending punctuation). We also sample paragraphs of languages uniformly. We continue training XLM-R on naturally occurring newline symbols for 200k training steps using a batch size of 512. We use a linearly increasing learning rate warmup from 0 to 1e-4, and decay the learning rate to 0 for the remaining 195k steps. We use the AdamW optimizer (Kingma and Ba, 2015). For the auxiliary objective as introduced by Minixhofer et al. (2023), we set the removal probability $p=0.25$ using the union of the 30 most common punctuation characters in every language. We then take the corresponding tokens as used by XLM-R as labels for the auxiliary objective.
For models without limited lookahead (cf. Table 9), we follow Minixhofer et al. (2023) and use a threshold of $0.01$ for sentence boundary detection. When using limited lookahead, we observe that the optimal threshold increases. We thus use a constant threshold of $0.025$ with limited lookahead.
Training of SaT+SM.
In our supervised mixture stage, we continue training SaT using the same context window of $256$. The training data now consists of sentence-segmented text, and we train SaT+SM predicting sentence-ending tokens.
For each language, we corrupt the data in two ways. In the first, we lowercase and remove all punctuation tokens. This aims to roughly emulate the output of an automatic speech recognition (ASR) system. In the second, we lowercase all text with probability $0.5$, remove all punctuation with probability $0.5$, and duplicate punctuation (e.g., changing ! to !!!) with geometric distribution scaling with the number of duplications (i.e., doubling with probability $0.5$, tripling with probability $0.25$, etc.). This aims to emulate user-generated text.
We generally pack sentences into chunks. For the uncorrupted sentences and sentences corrupted with the first scheme, we pack until each chunk fully fills up the model’s context window. For our second corruption scheme, we include $s$ sentences in each block, where $s$ is drawn from the same geometric distribution as used before.
We train with a batch size of 128, linear learning rate warmup from 0 to 3e-5 for 500 steps, and linearly decay for another 19,500 steps. We uniformly sample batches of sentences from a single language and evenly sample batches corrupted with one of the three corruption schemes. During inference, we use a constant threshold of $0.25$.
Training of SaT+LoRA.
For adaptation via LoRA, we use a learning rate of 3e-4. We train LoRA modules with AdamW for a target domain for 30 epochs, where we linearly warm up the learning rate for the first 10% of training, followed by a decay to 0 for the remaining training steps. We do not apply early stopping. We apply LoRA to the query and value matrices of the attention block, as well as the intermediate layer of the Transformer, using a rank $r=16$ and scaling factor $a=32$. We noticed this to positively impact performance at a comparably low computational cost. Moreover, similarly to WtPT and WtPPUNCT, we additionally tune the classification threshold on the same training data if more than $512$ sentences are available. We noticed that this helps performance in such cases. For verse segmentation, we use SaT models without limited lookahead, since, with verses, it is both helpful and desirable to rely on future verses.
LLM details.
We use default hyperparameters for both Llama 38B and Command R. For Command R, We used the Cohere API. This led to some API refusals, particularly for lyrics. We thus only consider chunks that were not refused when calculating metrics. To align input and output chunks using the Needleman-Wunsch sequence alignment algorithm, we use a gap penalty of $-0.5$, a gap extension penalty of $-0.5$, a match reward of $1$, and a mismatch penalty of $-0.5$. If no alignment is found, the LLM produced output that strongly deviated from the input chunk. We thus assign no sentence boundaries to the predictions of the LLMs for this input chunk.
We experiment with several prompts, optimizing performance on the training set, resulting in the following final prompt:
General LLM Prompt Separate the following text into sentences by adding a newline between each sentence. Do not modify the text in any way and keep the exact ordering of words! If you modify it, remove or add anything, you get fined \$1000 per word. Provide a concise answer without any introduction. Indicate sentence boundaries only via a single newline, no more than this!
[…]
For few-shot prompting, we append the prompt with When provided with multiple examples, you are to respond only to the last one. In addition, we indicate chunk $n$ with Input N: and Output N: , respectively.
Since it is a highly distinct task, we use the following prompt for verse segmentation:
LLM Lyrics Prompt Separate the following song’s lyrics into semantic units (e.g., verse, chorus, bridge, intro/outro, etc - similar to how they are presented in a lyrics booklet) via double newlines, but do not annotate them. Only include the song in the output, no annotations. Do not modify the song in any way and keep the exact ordering of words! If you modify it, remove or add anything, you get fined \$1000 per word. Indicate semantic units by double newlines.
A.3 Qualitative Examples
ASR output.
We show predictions of SaT+SM on parts of transcribed TED talks in different languages in Table 16.
Code-switching.
We also show predictions of SaT+SM on code-switching text in four language pairs in Table 17.
Verse segmentation.
In addition, we show predictions of SaT+LoRA on verse segmentation in Tables 18,  19, and  20 for songs of high, mid, and low levels of repetitiveness, respectively.
A.4 Additional Results
(a) Ersatz
(b) OPUS100
(c) UD
Figure 6: Avg. F1 vs. number of sentences used for adaptation, averaged over languages within a given dataset.
Results in English.
We provide an overview of the performance of different models on different corpora in Table 13.
Model OPUS100 UD Ersatz Macro Avg.
SpaCyM 88.8 91.7 94.0 91.5
Ersatz 87.6 89.1 97.5 91.4
Llama 38B 92.8 94.8 98.2 95.3
Command R 89.5 77.1 87.2 84.6
SaT 90.4 93.9 96.7 93.7
SaT+SM 94.6 96.7 98.3 96.5
NLTK 88.2 90.8 97.7 92.2
PySBD 59.6 75.3 73.9 69.6
SpaCyDP 89.0 91.3 98.5 92.9
WtP 90.6 94.5 96.5 93.9
WtPT 89.4 94.5 96.7 93.5
WtPPUNCT 94.7 96.9 98.6 96.7
SaT+LoRA 94.8 96.8 98.7 96.8
Table 13: English (en) sentence segmentation F1 scores. We use 3-layer versions of each model. Numerically best results are in bold, statistically indistinguishable ones from this best are underlined.
Legal data.
Table 21 shows the sentence segmentation performance of different models on MultiLegalSBD. Furthermore, Table 22 shows performance on MultiLegalSBD when applying the same corruptions as on Speeches, removing all casing and punctuation tokens.
More few-shot results.
Figure 6 shows the per-dataset few-shot domain adaptation results, comparing WtPT, WtPPUNCT, and SaT+LoRA.
Effect of stride.
For both SaT and WtP, we use a default stride of $64$ during evaluation. Since SaT operates on subwords but WtP on characters, this results in different scaling behaviors, illustrated in Figure 7.
Complete verse segmentation results.
We provide complete per-genre results for verse segmentation in Tables  23 and  24. Furthermore, Tables 25 and 26. show verse segmentation performance when applying the same corruptions as on Speeches, removing all casing and punctuation tokens.
Complete results on clean data.
Results of SaT, its variations, and other methods on all languages are shown in Tables 27-32.
Complete code-switching result.
We show results using both language codes on code-switched text for models relying on language codes in Table 14.
Figure 7: Sentence segmentation F1 scores vs. execution time across different strides (default 64), evaluated on Ersatz. We use the standard 3-layer variants of each model. Higher stride values result in faster inference.
Model es en de en tr de vi en Macro Avg.
NLTK 0.0/0.0 0.0/1.1 $0$/0.0 0.3/-
PySBD 0.0/0.0 2.1/2.1 -/0.0 -/0.0 0.5/0.5
SpaCyDP 0.0/17.6 8.6/8.0 -/12.2 -/11.3 12.2/-
WtP 20.8/38.6 39.0/31.4 33.5/21.0 22.7/25.5 29.1/29.0
WtPT 46.9/52.2 45.7/39.1 33.3/34.4 46.7/36.7 40.6/43.2
WtPPUNCT 60.7/62.1 60.1/58.1 39.9/41.0 59.0/50.7 53.0/54.9
Table 14: Complete sentence segmentation F1 scores for code-switched text for systems relying on language codes, where the first number corresponds to the first language shown.
Language iso Space UD OPUS100 Ersatz Speeches
Afrikaans af AfriBooms (425) 1.9k 1.3k
Amharic am 2.0k 514
Arabic ar PADT (680) 2.0k 1.5k 10.0k
Azerbaijani az 2.0k 6.8k
Belarusian be HSE (1.1k) 2.0k 6.2k
Bulgarian bg BTB (1.1k) 2.0k 10.0k
Bengali bn BRU (56) 2.0k 5.2k
Catalan ca AnCora (1.8k) 2.0k 10.0k
Cebuano ceb GJA (188) 55
Czech cs PDT (10.1k) 2.0k 1.7k 10.0k
Welsh cy CCG (953) 1.8k -
Danish da DDT (565) 2.0k 10.0k
German de GSD (977) 1.9k 2.0k 10.0k
Greek el GDT (456) 2.0k 10.0k
English en GUM (1.1k) 2.0k 7.7k 10.0k
Esperanto eo 2.0k 10.0k
Spanish es AnCora (1.7k) 2.0k 3.1k 10.0k
Estonian et EDT (3.2k) 2.0k 2.0k 5.9k
Basque eu BDT (1.8k) 2.0k 10.0k
Persian fa PerDT (1.5k) 2.0k 10.0k
Finnish fi TDT (1.6k) 2.0k 2.0k 10.0k
French fr GSD (416) 2.0k 1.7k -
Western Frisian fy 1.9k 46
Irish ga IDT (454) 2.0k -
Scottish Gaelic gd ARCOSG (545) 1.1k 10.0k
Galician gl TreeGal (400) 2.0k 7.9k
Gujarati gu 1.9k 1.0k 13
Hausa ha 2.0k 10.0k
Hebrew he IAHLTwiki (393) 2.0k 10.0k
Hindi hi HDTB (1.7k) 2.0k 2.5k 10.0k
Hungarian hu Szeged (449) 2.0k 10.0k
Armenian hy BSUT (595) 7.0k 104
Indonesian id PUD (1.0k) 2.0k 1.6k
Igbo ig 1.7k 10.0k
Icelandic is IcePaHC (5.2k) 2.0k 10.0k
Italian it ISDT (482) 2.0k 5.8k
Japanese ja GSD (543) 2.0k 1.1k 533
Javanese jv CSUI (125) 1.1k
Georgian ka 2.0k 10.0k
Kazakh kk KTB (1.0k) 1.9k 1.0k 4.1k
Khmer km 1.9k 2.4k 12
Kannada kn 906 10.0k
Korean ko Kaist (2.3k) 2.0k 215
Language iso Space UD OPUS100 Ersatz Speeches
Kurdish ku 1.9k 10.0k
Kirghiz ky 1.7k 3.0k
Latin la ITTB (2.1k) 10.0k
Lithuanian lt ALKSNIS (684) 2.0k 1.0k 2.1k
Latvian lv LVTB (2.3k) 2.0k 2.0k 10.0k
Malagasy mg 2.0k 103
Macedonian mk 2.0k 10.0k
Malayalam ml 2.0k 2.1k
Mongolian mn 4.2k 5.5k
Marathi mr UFAL (47) 2.0k 10.0k
Malay ms 1.9k 2.1k
Maltese mt MUDT (518) 2.0k 10.0k
Burmese my 2.0k 5.5k
Nepalese ne 1.9k 6.7k
Dutch nl Alpino (596) 2.0k 10.0k
Norwegian no Bokmaal (1.9k) 2.0k 2.1k
Panjabi pa 2.0k 3.8k
Polish pl PDB (2.2k) 2.0k 1.0k 548
Pushto ps 1.8k 2.7k 10.0k
Portuguese pt Bosque (1.2k) 2.0k 5.5k
Romanian ro Nonstandard (1.1k) 2.0k 2.0k 10.0k
Russian ru Taiga (881) 2.0k 991 10.0k
Sinhala si 2.0k 516
Slovak sk SNK (1.1k) 2.0k 10.0k
Slovenian sl SSJ (1.3k) 2.0k 10.0k
Albanian sq TSA (60) 2.0k 10.0k
Serbian sr SET (520) 2.0k 5.6k
Swedish sv LinES (1.0k) 2.0k 2.6k
Tamil ta TTB (120) 2.0k 1.0k 3.0k
Telugu te 2.0k 303
Tajik tg 2.0k 10.0k
Thai th PUD (1.0k) 2.0k 7.8k
Turkish tr IMST (983) 2.0k 3.0k 3.8k
Ukrainian uk IU (892) 2.0k 10.0k
Urdu ur UDTB (535) 1.9k 7.8k
Uzbek uz 2.0k 3.8k
Vietnamese vi VTB (800) 1.9k 10.0k
Xhosa xh 1.9k 5.6k
Yiddish yi 1.3k 2.6k
Yoruba yo YTB (318) 9.4k 10.0k
Chinese zh GSDSimp (500) 2.0k 2.0k 1.7k
Zulu zu 1.9k 8.1k
Table 15: List of the 85 languages considered, whether they generally use whitespace to split sentences, and the corresponding evaluation dataset size, measured in sentences. For UD, we use UDv2.13, where the treebank name used is also shown. We use Speeches only in pairwise evaluations.
English (en) we use science to create something wonderful
we use story and artistic touch to get us to a place of wonder
this guy wall-e is a great example of that
he finds beauty in the simplest things
but when he came in to lighting we knew we had a big problem
we got so geeked-out on making wall-e this convincing robot that we made his binoculars practically optically
     perfect
laughter
(*) his binoculars are one of the most critical acting devices he has
he doesn’t have a face or even traditional dialogue for that matter
so the animators were heavily dependent on the binoculars to sell his acting and emotions
we started lighting and we realized the triple lenses inside his binoculars were a mess of reflections
he was starting to look glassy-eyed
German (de) aber ich habe auch das marfan-syndrom
das ist eine erbkrankheit
1992 nahm ich an einer genetikstudie teil
zu meinem entsetzen erfuhr ich dass wie sie hier sehen meine aorta ascendens nicht im normalbereich war die
      grüne linie hier unten
alle hier im raum werden bei 3,2 und 3,6 cm liegen
ich war bereits bei 4,4
wie sie sehen können erweiterte sich meine aorta zunehmend und allmählich geriet ich an den punkt dass eine
      operation nötig sein würde
die angebotene operation war ziemlich gruselig
anästhesie öffnen des brustkorbs man hängt sie an eine künstliche herzlungenmaschine lässt ihre
      körpertemperatur auf etwa 18 grad fallen hält ihr herz an schneidet die aorta raus ersetzt sie mit einer klappe und aorta aus plastik
und am wichtigsten verdonnert sie lebenslang zu antikoagulationstherapie
(*) normalerweise mit warfarin
der gedanke an diese operation war nicht gerade ansprechend
French (fr) j’ai montré mon intro et j’ai mis la scène de la méduse
le réalisateur est resté silencieux pendant un très long moment (|) assez long pour que je puisse me dire oh non
      c’est foutu
et il a commencé à applaudir
puis le concepteur artistique
(*) et finalement toute la salle
c’est pour ces moments que je fais ce travail
le moment où tout fait sens et où l’on crée un monde auquel on peut croire
on utilise la science et la programmation pour créer ces mondes incroyables
on utilise les histoires et l’art pour leur donner vie
c’est la coexistence de l’art et la science qui transforme le monde en un lieu magique un lieu avec une âme un
      lieu auquel on peut croire un lieu où les choses qu’on imagine deviennent réelles – et un monde où tout d’un
      coup une fille réalise qu’elle n’est pas seulement une scientifique mais aussi une artiste
merci (*) applaudissements
Italian (it) cosa state leggendo
beau lotto
(*) cosa state leggendo
mancano metà delle lettere
giusto
non c’è nessuna ragione a priori perché una h debba comparire tra la w e la a
ma ne collocate una lì
perché
perché nella statistica della vostra esperienza passata sarebbe stato utile fare così
quindi lo fate di nuovo
e tuttavia non collocate una lettera dopo quella prima t
perché (|) perché non si sarebbe dimostrato utile nel passato
quindi non lo fate di nuovo
Table 16: Examples of predictions of SaT+SM taken from random positions from transcribed TED talks in four langauges. (|) marks a missing sentence boundary (false negative), and (*) marks a wrongly inserted sentence boundary (false positive). All others are correctly segmented, according to the ground truth segmentation.
German- English (de-en) its about the rundfunkstaatsvertrag and the licence you need to stream to more than 500 viewers
just go to the amt on your day off arrive at 9 and bring all the papers and a book
hat ihr kein editor angekreidet
guess op thought everyone pays freiwillige gesetzliche krankenversicherung und pflegeversicherung which
      is around 780eurs month per person family depends (|) is the date that matters
the bescheinigungszeitraum on my ausdruck der elektronischen lohnsteuerbescheinigung für 2013 or my
      first anmeldung
na anti-establishment anti-kapitalismus oder generell anti-zwang
(*) zum beispiel (|) ich bin software engineer mit data-management data-security background und er is
      front-end mobile dev
noch besser in flughäfen in england kommt keine höflich message wie achten sie bitte auf ihr eigenes
      gepäck… sondern direkt eine drohung for security reasons baggage left unattended will be removed and
      destroyed
da bleibt irgendwie nicht mehr viel übrig
Spanish- English (es-en) in the morning over there cada vez que yo decía algo él me decía algo
the best thing about her ella no complain you know (|) tiene she has a great personality
hasta que tú pushed the wrong button
linda lópez la testing
ay but she’s cool
el teniente y ella han tenido tú sabes conflicts
entonces tina es the computer person
so ella ella es la jefa de linda
sí i i have a room
no
(*) pero tal vez consigue un roommate
un roommate (|) mandó un e-mail diciendo que le había que había otra persona en la dirección en lugar de
      ella
yo me dio a entender según como leí yo que era ella e edith
Turkish- German (tr-de) ja bence ich probiere es einfach
ja vor allem sınav olduğu zaman muss ja muss ja schon so sein
geçen sene ich weiß noch eh sınavların olduğu günlerde (|) ja tam denk geldi weißt du (|) die woche noch
      ich denke mir so tutayım mı tutmayayım mı kann gar nicht mehr
ben tutmuştum bir tanesinde
(*) und ich dachte so tamam bittin sen die
(*) hani das ist nicht gut gegangen
(*) die prüfung das war auch so
(*) ich konnte mich gar konzentrieren überhaupt nicht
hani yemek de değil
(*) weißt du einfach nur wasser
ja bir de o geçen sene da war es so heiß
Vietnamese- English (vi-en) bởi vậy lux đường có overconfident (—) ai cũng cần phải improve (—) nên lux phải phải phải đưa cho chị ti với
      anh alex check nha (—) thì design đến đâu rồi lux (—) cái graphic design của lux
(*) lux design được đến đâu rồi come up with idea để ghi ra
(*) chị ti sẽ cùng help lux to write down
(*) hoặc là ở woden qua đây ăn dinner với chị ti
get it out of the way là done với một cái đó rồi là xong
thái writing cũng đâu có good đâu
deadline như vậy được chưa
vậy là tối mai lux biết spend time write it tomorrow (—) được rồi ta design như vậy nè
mọi thứ là lux phải plan ahead như vậy chứ
Table 17: Examples of predictions of SaT+SM taken from random positions from code-switching text in four language pairs. (|) marks a missing sentence boundary (false negative), and (*) marks a wrongly inserted sentence boundary (false positive). All others are correctly segmented according to the ground truth segmentation. There are many ambiguous sentence boundaries in these corpora.
Original (non-corrupted) song Corrupted song
Have yourself a merry little Christmas Let your hearts be light From now on Our troubles will be out of sight Have yourself a merry little Christmas Make the Yule-tide gay From now on Our troubles will be miles away. Here we are as in olden days Happy golden days of yore Faithful friends who are dear to us Will be near to us once more Through the years We all will be together If the Fates allow Until then we’ll have to muttle through somehow So have yourself a merry little Christmas now Here we are as in olden days Happy golden days of yore Faithful friends who are dear to us Will be near to us once more Through the years We all will be together If the Fates allow Until then we’ll have to muttle through somehow So have yourself a merry little Christmas now Have yourself a merry little Christmas now Merry Christmas i’d never leave the perfect girl or rip apart the perfect world just up and leave in the middle of a song i’d never pack my things in a silverado drive on out to colorado just to find some freedom i thought was gone (|) ooooh there are things i’d never do but here i am in this hotel room thinking bout you and what i ’ve done oh what have i done head in my hands thinking about a lot of things i wish that i could change it’s sad but it’s true i ’ve done a lot of things i’d never do i’d never ever work so much that i’d lose sight i’d lose touch of everything a man could ever want i’d never lose my cool and say those words that cut just like a blade and leave you dying crying all alone i’d never leave the perfect girl or rip apart the perfect world just up and leave in the middle of a song
Table 18: Examples of predictions of SaT+LoRA taken from songs categorized as Country (High Repetitiveness).
(|) marks a missing verse boundary (false negative), and (*) marks a wrongly inserted verse boundary (false positive). All others are correctly segmented according to the ground truth segmentation. While the task was only to segment songs into verses and no line segmentation was provided, we format songs using both lines and verses for clarity.
Original (non-corrupted) song Corrupted song
Hope, a new beginning Time, time to start living Just like just before we died There’s no going back to the place we started from Hurt, falling through fingers Trust, trust in the feeling There’s something left inside There’s no going back to the place we started from All secrets known Calm, old wounds are healing Strong, truth is worth saving I want to feel alive There’s no going back to the place we started from All secrets known lock me up inside my room leave me without toys and food keep that monster in my bed just remember i’m not dead you forget my memory lives way beyond these walls you forget my indecision’s taking all control once in a far land i grabbed you and you woke me up to my origin people have to understand my innocence has gone (|) go beyond my urge or make an effort to living on my own plagued by images once in a far land i grabbed you and you woke me up to my origin
Table 19: Examples of predictions of SaT+LoRA from songs categorized as Alternative Metal (Mid Repetitiveness).
Original (non-corrupted) song Corrupted song
Let me chirp these fools Juice got weed Juice got pills Juice got the work on the corner cutting deals Juice know you haters out there snitching ain’t for real So Juice got some gang niggas down for the kill Juice know the feds got surveillance on the field We never had a job but we sitting on a mill We ball out in the club wit our niggas staying trill We never wrote a check just them big face bills A player drinking Makers Marka, cranberry vodka Wearing a mink coat thats furry as Chewbacca I saw ya main girl and a player had to stop her Her name wasn’t Silkk but her face was The Shocker The feds taking pictures of us balling but I got ’em A 7 footer hole for his body we gonna drop ’em We always on the grind we be watching when they watching And when they turn they back its the clucka-clucka-rock ’em yeah! If you boys got beef we can (roll wit it) In the club or the street we can (go wit it) It don’t make me none (blow for blow wit it) Crack his head wit a gun (I’ma sho split it) We got them tones in the club and them bulletproof vests Them three fifty seven titanium Smith-N-Wess And plus we deep as hell and prepared to bust You gonna have hell if you fuck wit us and thats whats up (*) The whole club we maintain These hydrashock bullets mushroom in ya brain We in bed with the med we give ’em something to do Cause clown ass niggas love to act a fool My hood is real nigga my hood ain’t fake My hood is home nigga everything straight My hood will rob you with mask on they face My hood will do it to put food on they plate My hood ain’t tame dog they wanna jump fool My hood they hang together they all jump you And if you don’t believe me then come to my hood And you will see that it ain’t all good zaytoven on the track (|) zay-tiggy gucci gucci so watch entertainment lets go they call me chef-boy-r.g. but hold that thought its a kodak moment but hold that thought hurricane wrist game turn that junk off hot as piggly wiggly cant kermit the frog dog early in the mornin i aint even yawnin cookin up a cake like i’m doin a performance when it come to flossin i aint even talkin diamonds on my joint got my chevy moonwalkin 10 bricks on my bart simpson just look my watch 35 pounds of kush my ring 36 oz’s my nig my bracelet 500 lbs of mid a gucci wrapped tour bus yall hoes follow us party pack pills man hoes gonna swallow us naturally a loner but love my kid mix the soda with the cola i can buy me a friend new swag somethin like trap house times 10 ery nigga round me bust heads, ya-dig iced out grill i can’t buy that bullshit i’m wit some street shit, like a reverend in the pulpit they call me chef-boy-r.g. but hold that thought its a kodak moment but hold that thought hurricane wrist game turn that junk off hot as piggly wiggly cant kermit the frog dog
Table 20: Examples of predictions of SaT+LoRA from songs categorized as Southern Hip Hop (Low Repetitiveness).
Model fr es it en de Macro Avg.
Judg. Laws Judg. Laws Judg. Laws Judg. Judg. Laws
SpaCyM 82.7 61.2 67.7 85.2 73.6 53.8 81.4 65.8 67.9 71.0
Ersatz 81.5 51.7 63.6 81.9 - - 79.5 59.8 68.5 69.5
Llama 38B 85.5 52.7 70.4 66.5 81.1 77.2 89.2 78.3 83.4 76.0
Command R 62.5 51.2 55.5 52.8 56.2 70.3 69.0 55.7 64.3 59.7
SaT 82.7 86.8 68.8 73.2 85.1 71.2 82.6 67.8 86.4 78.3
SaT+SM 85.1 95.8 80.1 89.3 88.3 80.6 93.1 81.3 92.7 87.4
NLTK 75.6 51.5 65.2 87.7 72.9 45.3 76.3 65.1 73.3 68.1
PySBD 74.2 50.5 60.8 79.7 74.1 55.0 75.0 67.6 70.4 67.5
SpaCyDP 71.6 74.3 64.9 87.3 74.0 54.4 84.5 68.2 65.2 71.6
WtP 87.8 63.0 75.6 84.6 84.3 80.0 88.8 79.7 85.8 81.1
WtPT 87.0 80.7 76.3 87.1 85.0 79.5 88.7 80.4 85.9 83.4
WtPPUNCT 96.8 98.4 88.7 94.6 94.0 96.9 96.3 87.3 93.7 94.1
mLSBD-T Mono Specific 96.4 97.7 88.7 93.8 93.7 98.0 95.3 86.9 93.5 93.8
mLSBD-T Mono Both 96.5 98.5 88.2 93.7 87.1 84.4 95.3 87.6 92.7 91.6
mLSBD-T Multi Specific 96.4 98.9 89.0 94.8 94.5 98.1 95.7 87.5 93.6 94.3
mLSBD-T Multi Both 96.6 98.9 88.7 94.8 94.6 98.2 95.6 78.6 93.2 93.2
SaT+LoRA @10 95.1 96.6 86.5 93.5 94.0 85.6 96.3 87.7 97.3 92.5
SaT+LoRA @100 96.7 97.0 89.3 94.0 95.4 87.1 97.0 88.8 97.3 93.6
SaT+LoRA 96.8 98.5 89.1 94.4 95.5 98.3 97.3 88.9 97.1 95.1
Table 21: Sentence segmentation F1 score for legal data (MultiLegalSBD). We take the macro F1 scores over documents within a given category. @N correspond to using a maximum of N documents per category for adaptation, respectively. Transformer-based mLSBD-T baselines are taken from Brugger et al. (2023). For these domain-specific baselines, mono and multi correspond to models trained on documents from only one or all languages, respectively. Both corresponds to models trained on both laws and judgments, whereas specific corresponds to models trained on a given category (laws/judgments). Best per-category results are in bold.
Model fr es it en de Macro Avg.
Judg. Laws Judg. Laws Judg. Laws Judg. Judg. Laws
SpaCyM 0.0 0.0 1.5 0.0 0.4 0.0 0.2 2.2 0.0 0.5
Ersatz 0.7 0.1 4.7 0.0 - - 0.6 4.1 - 1.7
Llama 38B 60.0 40.7 50.5 55.9 54.2 35.6 69.2 59.2 74.2 55.5
Command R 54.3 57.7 42.8 44.5 46.4 42.9 56.0 45.4 64.1 50.5
SaT 63.7 80.6 54.9 63.3 56.0 55.9 49.9 57.6 75.0 61.9
SaT+SM 68.8 89.5 65.3 83.3 62.3 71.9 74.5 72.0 80.9 74.3
NLTK 1.4 0.0 3.8 0.0 1.3 1.6 0.2 4.9 0.0 1.5
PySBD 21.7 0.2 3.8 0.0 20.0 0.2 1.3 0.4 0.0 5.3
SpaCyDP 7.9 4.4 4.6 0.0 5.3 2.3 2.4 13.1 8.8 5.4
WtP 32.2 19.5 38.0 41.1 36.9 28.8 28.1 25.0 19.0 29.8
WtPT 48.9 64.7 51.8 71.5 46.5 43.8 54.5 52.4 62.0 55.1
WtPPUNCT 65.0 83.4 67.2 83.1 58.3 66.9 73.5 73.0 84.1 72.7
mLSBD-T Mono Specific 9.6 45.1 7.9 13.2 6.9 47.1 3.6 12.1 0.3 16.2
mLSBD-T Mono Both 9.0 43.0 8.1 15.6 6.7 39.6 3.6 7.7 0.6 14.9
mLSBD-T Multi Specific 8.8 44.4 7.3 34.4 6.9 47.2 2.5 9.2 0.7 17.9
mLSBD-T Multi Both 8.9 43.5 5.6 24.6 6.9 47.4 1.3 2.4 0.4 15.7
SaT+LoRA @10 70.8 82.9 67.8 79.9 63.5 62.6 80.6 78.6 93.5 75.6
SaT+LoRA @100 76.9 86.6 75.9 84.1 69.3 75.3 84.2 84.1 93.5 81.1
SaT+LoRA 77.5 90.2 76.1 85.5 71.2 87.5 84.2 83.8 93.5 83.3
Table 22: Sentence segmentation F1 score for corrupted legal data (MultiLegalSBD), where we remove all casing and punctuation tokens. We take the macro F1 scores over documents within a given category.
Repetitiveness
High Low
Model Country Punk Rock Pop Punk Southern Hip Hop Gangsta Rap
$\text{SSM}_{\text{string}}$
Llama 38B 47.0 50.2 49.5 34.7 32.8
Command R 35.3 39.7 39.2 27.5 29.9
WtPPUNCT @100 56.5 56.1 54.9 43.3 42.4
WtPPUNCT @1000 58.9 58.8 57.5 45.9 43.9
WtPPUNCT 58.9 58.8 57.5 45.9 43.9
SaT+LoRA @100 66.7 67.5 69.2 53.0 50.3
SaT+LoRA @1000 76.7 76.0 76.8 64.5 60.9
SaT+LoRA 79.3 76.8 77.6 64.2 60.3
Table 23: Complete verse segmentation F1 scores for songs categorized as low and high repetitiveness. Categorization of songs into genres and repetitiveness taken from Fell et al. (2018). We report the macro average over songs. Values for $\text{SSM}_{\text{string}}$ taken from Fell et al. (2018), with lyrics already pre-segmented into lines. @N corresponds to using a maximum of N and 1000 songs per genre for adaptation, respectively.
Mid Repetitiveness
Model Rock Pop RnB Soul Alt. Rock Alt. Metal Indie Rock Pop Rock Hard Rock
$\text{SSM}_{\text{string}}$ 64.8 66.6 65.6 63.0 67.9 68.5 65.6 65.8 67.7
Llama 38B 48.2 47.0 45.7 48.7 49.3 47.6 48.4 47.1 48.7
Command R 37.8 36.3 33.3 36.5 40.1 39.1 40.2 37.6 38.4
WtPPUNCT @100 57.5 55.9 51.4 52.5 59.9 58.6 56.3 57.5 57.0
WtPPUNCT @1000 60.5 57.7 53.1 55.0 63.0 60.5 59.8 58.1 59.4
WtPPUNCT 60.7 58.2 53.5 55.0 63.3 60.5 60.1 58.3 59.5
SaT+LoRA @100 65.4 63.8 61.5 59.8 64.9 68.6 63.9 63.8 64.1
SaT+LoRA @1000 74.7 72.7 71.2 71.1 75.3 77.7 72.6 74.5 75.7
SaT+LoRA 78.1 75.6 73.4 71.7 77.4 77.7 73.5 75.6 76.6
Table 24: Complete verse segmentation F1 scores for songs categorized as mid repetitiveness.
Repetitiveness
High Low
Model Country Punk Rock Pop Punk Southern Hip Hop Gangsta Rap
$\text{SSM}_{\text{string}}$ 70.2 70.9 72.7 47.0 47.7
Llama 38B 54.6 53.5 57.1 36.6 36.2
Command R 38.8 42.6 41.4 35.0 26.3
WtPPUNCT @100 48.9 52.5 50.0 35.5 36.4
WtPPUNCT @1000 51.7 55.6 53.1 36.5 39.5
WtPPUNCT 51.9 55.6 53.1 36.5 39.5
SaT+LoRA @100 66.7 67.2 67.7 44.9 48.4
SaT+LoRA @1000 72.7 71.8 74.5 54.7 55.8
SaT+LoRA 75.2 71.4 74.7 54.8 55.8
Table 25: Complete verse segmentation F1 scores for songs categorized as low and high repetitiveness, where we remove all casing and punctuation tokens. We report the macro average over songs.
Mid Repetitiveness
Model Rock Pop RnB Soul Alt. Rock Alt. Metal Indie Rock Pop Rock Hard Rock
$\text{SSM}_{\text{string}}$
Llama 38B 53.8 51.7 47.4 51.9 54.1 52.3 54.0 52.8 52.2
Command R 41.1 37.5 34.0 40.0 41.4 40.7 42.6 39.5 40.6
WtPPUNCT @100 51.5 47.0 41.5 47.3 52.6 52.7 52.4 48.8 50.4
WtPPUNCT @1000 52.8 48.9 43.1 48.6 54.9 54.2 55.1 51.2 52.0
WtPPUNCT 53.1 49.4 43.2 48.6 55.0 54.2 54.8 51.0 52.6
SaT+LoRA @100 64.5 62.2 58.8 60.8 67.4 66.9 62.7 63.5 63.6
SaT+LoRA @1000 70.4 68.5 65.2 65.6 71.4 73.1 69.9 68.3 71.4
SaT+LoRA 73.2 71.5 67.1 65.1 73.3 73.5 69.5 70.4 72.7
Table 26: Complete verse segmentation F1 scores for songs categorized as mid repetitiveness, where we remove all casing and punctuation tokens. We report the macro average over songs.
Model af am ar az be bg bn ca ceb cs cy da de el en
UD SpaCyM 98.3 - 79.1 - 80.0 93.8 47.9 98.8 98.5 89.6 98.9 94.9 87.7 93.8 91.7
Ersatz - - 79.7 - - - - - - 89.3 - - 92.4 - 89.1
Llama 38B 100.0 - 80.1 - 88.1 97.3 96.9 99.4 99.7 93.0 98.7 95.2 96.7 94.7 94.8
Command R 86.1 - 75.6 - 68.1 76.9 59.2 76.6 89.9 68.9 80.8 73.8 85.3 66.7 77.1
SaT 99.1 - 82.9 - 89.0 97.9 96.2 98.9 99.7 91.5 99.1 95.5 96.4 97.5 93.9
SaT+SM 100.0 - 84.5 - 93.5 99.3 100.0 99.7 99.7 94.3 99.4 98.5 97.8 97.9 96.7
NLTK - - - - - - - - - 89.1 - 94.4 92.6 92.7 90.8
PySBD - - 28.1 - - 74.9 - - - - - 72.6 80.0 91.0 75.3
SpaCyDP - - - - - - - 99.8 - - - 94.0 96.7 94.0 91.3
WtP 98.3 - 80.5 - 88.9 98.2 93.5 98.3 99.7 92.3 99.2 95.1 95.6 97.3 94.5
WtPT 99.0 - 86.4 - 88.8 98.1 - 98.4 - 92.0 99.2 94.3 95.8 97.7 94.5
WtPPUNCT 99.9 - 87.4 - 91.9 99.6 - 99.6 - 95.4 99.5 98.9 96.5 97.8 96.9
SaT+LoRA 100.0 - 86.6 - 91.2 99.4 - 99.9 - 95.2 99.6 98.7 96.8 98.9 96.8
OPUS100 SpaCyM 41.9 6.3 57.9 72.3 33.9 93.4 37.2 88.0 - 87.3 25.8 90.2 72.9 89.2 88.8
Ersatz - - 59.2 - - - - - - 86.5 - - 73.1 - 87.6
Llama 38B 65.4 36.2 62.2 76.8 54.5 94.6 80.0 90.9 - 91.3 46.3 93.2 83.8 94.3 92.8
Command R 55.8 6.6 44.5 60.5 36.1 57.9 4.8 66.0 - 69.0 39.8 75.4 76.2 65.1 89.5
SaT 78.4 58.0 67.0 75.0 70.4 93.5 80.5 87.7 - 89.2 72.0 90.9 78.2 92.0 90.4
SaT+SM 86.2 70.8 65.2 85.3 87.8 96.2 86.1 93.0 - 94.3 79.6 94.0 86.9 95.9 94.6
NLTK - - - - - - - - - 86.7 - 89.9 73.3 84.8 88.2
PySBD - 5.9 38.0 - - 72.9 - - - - - 70.2 66.5 62.7 59.6
SpaCyDP - - - - - - - 87.3 - - - 90.2 74.0 91.1 89.0
WtP 74.6 58.2 64.5 74.9 71.7 93.2 77.9 87.7 - 87.5 68.7 88.2 76.7 90.9 90.6
WtPT 76.4 63.7 64.6 74.6 72.5 92.8 82.1 88.6 - 90.0 74.2 90.1 84.6 91.9 89.4
WtPPUNCT 87.8 70.5 76.1 83.0 89.1 96.2 86.5 94.0 - 94.9 81.0 94.5 89.4 95.9 94.7
SaT+LoRA 88.5 75.9 79.1 85.0 89.4 96.4 88.6 94.6 - 95.0 83.0 95.2 90.1 96.1 94.8
Ersatz SpaCyM - - 91.1 - - - - - - 96.4 - - 93.5 - 94.0
Ersatz - - 92.8 - - - - - - 96.7 - - 95.4 - 97.5
Llama 38B - - 92.4 - - - - - - 95.9 - - 97.3 - 98.2
Command R - - 55.8 - - - - - - 66.4 - - 75.9 - 87.2
SaT - - 89.7 - - - - - - 94.3 - - 96.6 - 96.7
SaT+SM - - 92.3 - - - - - - 98.5 - - 97.2 - 98.3
NLTK - - - - - - - - - 96.7 - - 95.3 - 97.7
PySBD - - 46.2 - - - - - - - - - 95.3 - 73.9
SpaCyDP - - - - - - - - - - - - 96.3 - 98.5
WtP - - 87.0 - - - - - - 93.6 - - 95.3 - 96.5
WtPT - - 88.7 - - - - - - 93.9 - - 95.6 - 96.7
WtPPUNCT - - 92.7 - - - - - - 99.0 - - 99.3 - 98.6
SaT+LoRA - - 93.1 - - - - - - 98.5 - - 99.3 - 98.7
Table 27: Sentence segmentation test F1 scores on languages af-en. Results are shown using 3-layer variations of all models. Numerically best results are in bold, statistically indistinguishable ones from this best are underlined.
Model eo es et eu fa fi fr fy ga gd gl gu ha he hi
UD SpaCyM - 98.6 93.6 95.7 99.8 92.7 96.1 - 96.3 62.9 92.1 - - 93.9 -
Ersatz - 97.5 92.9 - - 92.8 97.3 - - - - - - - 99.5
Llama 38B - 98.5 94.1 97.3 99.7 95.9 99.1 - 94.5 67.6 94.6 - - 93.4 99.8
Command R - 81.8 74.9 79.4 95.5 75.3 92.3 - 73.5 54.2 73.7 - - 77.2 91.8
SaT - 97.0 92.8 97.0 98.5 93.8 97.5 - 87.3 68.1 97.4 - - 94.2 96.2
SaT+SM - 99.4 98.2 100.0 99.9 97.2 98.3 - 95.5 84.3 98.9 - - 95.5 99.9
NLTK - 98.5 93.6 - - 92.6 97.0 - - - - - - - -
PySBD - 46.2 - - 98.8 - 62.1 - - - - - - - 99.8
SpaCyDP - 99.1 - - - 94.9 91.6 - - - - - - - -
WtP - 96.5 92.6 97.1 96.6 92.1 96.4 - 84.3 71.2 97.5 - - 95.1 96.1
WtPT - 96.9 92.5 97.3 97.8 92.7 96.6 - 90.4 70.7 98.7 - - 96.1 96.8
WtPPUNCT - 99.7 98.0 99.9 100.0 98.1 98.3 - 98.2 79.6 98.6 - - 97.1 99.9
SaT+LoRA - 99.6 97.7 100.0 100.0 97.7 98.9 - 98.8 82.2 98.7 - - 96.4 99.9
OPUS100 SpaCyM 88.4 90.4 87.0 80.7 54.5 92.9 85.1 21.7 61.0 36.4 88.2 5.2 88.5 91.8 52.8
Ersatz - 90.0 87.4 - - 92.7 86.1 - - - - 21.2 - - 58.0
Llama 38B 91.6 93.6 89.0 84.5 60.4 93.9 91.4 42.4 71.3 62.8 90.8 33.0 88.1 91.6 59.6
Command R 84.6 80.1 67.2 57.8 42.9 68.4 80.8 30.6 55.9 50.5 71.6 16.7 65.8 62.7 39.3
SaT 90.3 91.1 84.6 82.2 56.8 91.3 87.3 64.9 82.3 81.6 87.9 71.5 83.8 89.4 62.9
SaT+SM 95.2 95.3 93.0 90.0 60.1 95.3 92.6 91.0 80.8 82.5 93.1 83.4 91.4 92.2 72.8
NLTK - 89.8 87.6 - - 93.1 85.8 - - - - - - - -
PySBD - 67.6 - - 41.3 - 80.9 - - - - - - - 23.0
SpaCyDP - 88.0 - - - 92.1 84.1 - - - - - - - -
WtP 90.9 89.9 82.5 84.4 59.6 90.8 86.8 44.4 77.8 83.5 88.5 69.2 82.8 90.2 65.1
WtPT 90.5 91.4 87.7 86.0 59.3 92.5 - 61.0 77.3 83.7 88.9 69.6 88.9 89.4 64.3
WtPPUNCT 95.4 95.0 94.6 91.6 72.7 95.5 - 88.1 87.5 92.7 93.9 76.5 91.5 93.8 76.3
SaT+LoRA 95.8 95.8 94.7 92.8 74.1 96.3 - 88.8 90.6 94.5 94.6 80.5 91.0 94.1 81.5
Ersatz SpaCyM - 97.2 97.0 - - 95.0 96.4 - - - - 3.8 - - 17.9
Ersatz - 96.6 98.0 - - 96.0 96.3 - - - - 94.3 - - 96.8
Llama 38B - 98.3 97.0 - - 96.7 97.9 - - - - 93.1 - - 97.3
Command R - 76.8 77.2 - - 78.8 84.6 - - - - 74.8 - - 84.9
SaT - 98.4 95.9 - - 97.6 97.3 - - - - 92.0 - - 95.5
SaT+SM - 99.5 98.9 - - 98.4 98.5 - - - - 70.7 - - 98.1
NLTK - 96.4 97.5 - - 95.9 96.0 - - - - - - - -
PySBD - 84.2 - - - - 96.0 - - - - - - - 87.5
SpaCyDP - - - - - - 91.0 - - - - - - - -
WtP - 98.7 96.0 - - 97.4 97.2 - - - - 89.7 - - 93.9
WtPT - - 95.7 - - 97.2 96.6 - - - - 88.9 - - 94.6
WtPPUNCT - - 98.0 - - 99.4 98.4 - - - - 96.7 - - 96.3
SaT+LoRA - - 99.0 - - 98.6 99.1 - - - - 95.4 - - 97.3
Table 28: Sentence segmentation test F1 scores on languages eo-hi.
Model hu hy id ig is it ja jv ka kk km kn ko ku ky
UD SpaCyM 98.3 11.3 97.8 - 95.1 92.3 0.0 97.9 - 96.1 - - 99.8 - -
Ersatz - - - - - - 93.4 - - 95.6 - - - - -
Llama 38B 98.1 92.0 99.3 - 94.1 98.1 97.9 99.1 - 98.5 - - 99.9 - -
Command R 73.2 50.3 91.7 - 74.6 88.8 90.7 83.3 - 86.1 - - 79.9 - -
SaT 97.0 97.6 98.5 - 73.4 96.3 96.1 98.8 - 95.8 - - 99.6 - -
SaT+SM 99.5 98.4 99.6 - 95.7 98.7 97.1 99.3 - 99.3 - - 100.0 - -
NLTK - - - - - 95.2 - - - - - - - - -
PySBD - 92.8 - - - 74.9 97.9 - - 95.5 - - - - -
SpaCyDP - - - - - 99.5 97.8 - - - - - 99.9 - -
WtP 96.0 96.0 98.0 - 85.8 93.7 93.4 97.8 - 97.4 - - 99.2 - -
WtPT 96.3 96.2 - - 88.9 93.7 95.7 - - 82.8 - - 99.4 - -
WtPPUNCT 99.5 98.5 - - 96.6 99.3 98.1 - - 84.7 - - 99.9 - -
SaT+LoRA 99.5 98.3 - - 96.8 99.7 98.0 - - 96.9 - - 99.9 - -
OPUS100 SpaCyM 93.0 24.2 89.6 29.8 95.0 85.8 0.1 - 38.3 42.1 0.0 9.8 50.9 26.8 21.7
Ersatz - - - - - - 28.8 - - 37.9 0.0 - - - -
Llama 38B 94.2 73.1 92.3 39.3 95.1 90.7 9.0 - 89.2 56.4 9.9 26.7 70.3 41.9 58.3
Command R 62.6 46.7 71.3 28.0 77.7 78.6 6.0 - 25.6 34.5 4.3 5.5 38.9 29.8 26.5
SaT 92.3 85.3 86.6 81.7 93.7 87.0 83.1 - 75.7 80.3 70.4 70.2 72.4 77.2 84.1
SaT+SM 95.8 90.6 93.7 92.1 96.5 90.9 78.0 - 93.6 92.2 86.6 87.9 78.9 91.1 91.8
NLTK - - - - - 87.5 - - - - - - - - -
PySBD - 58.4 - - - 70.2 43.4 - - 35.7 - - - - -
SpaCyDP - - - - - 85.3 42.6 - - - - - 46.6 - -
WtP 91.6 85.1 88.9 78.6 93.9 84.6 44.6 - 91.3 73.4 71.8 64.5 56.7 78.1 84.5
WtPT 92.1 - 89.5 82.1 94.4 88.4 79.5 - 91.1 74.9 71.1 60.4 70.4 66.4 84.4
WtPPUNCT 96.2 - 93.9 90.2 96.6 93.4 86.7 - 92.8 92.1 79.0 78.0 81.6 85.1 90.3
SaT+LoRA 96.3 - 94.1 92.3 96.6 94.3 90.5 - 93.3 92.1 87.8 83.9 81.8 90.9 92.0
Ersatz SpaCyM - - - - - - 0.0 - - 96.9 0.0 - - - -
Ersatz - - - - - - 85.7 - - 99.6 31.3 - - - -
Llama 38B - - - - - - 87.1 - - 99.0 85.6 - - - -
Command R - - - - - - 59.7 - - 86.7 6.7 - - - -
SaT - - - - - - 86.5 - - 97.1 89.4 - - - -
SaT+SM - - - - - - 89.1 - - 99.7 83.9 - - - -
NLTK - - - - - - - - - - - - - - -
PySBD - - - - - - 87.0 - - 64.7 - - - - -
SpaCyDP - - - - - - 91.0 - - - - - - - -
WtP - - - - - - 80.2 - - 96.3 70.2 - - - -
WtPT - - - - - - 81.5 - - 95.7 91.4 - - - -
WtPPUNCT - - - - - - 96.7 - - 99.7 92.0 - - - -
SaT+LoRA - - - - - - 94.6 - - 99.8 92.0 - - - -
Table 29: Sentence segmentation test F1 scores on languages hu-ky.
Model la lt lv mg mk ml mn mr ms mt my ne nl no pa
UD SpaCyM 0.0 93.8 98.5 - - - - 92.5 - 85.7 - - 92.8 96.9 -
Ersatz - 92.2 96.9 - - - - - - - - - - - -
Llama 38B 90.0 94.3 97.0 - - - - 90.0 - 94.4 - - 96.0 96.5 -
Command R 65.1 65.7 69.5 - - - - 71.6 - 70.1 - - 66.3 69.5 -
SaT 67.8 97.1 96.4 - - - - 83.5 - 86.7 - - 92.6 98.6 -
SaT+SM 96.6 97.9 99.2 - - - - 100.0 - 93.0 - - 95.7 99.0 -
NLTK - - - - - - - - - - - - 95.7 95.6 -
PySBD - - - - - - - 60.3 - - - - 93.6 - -
SpaCyDP - 92.0 - - - - - - - - - - 93.1 - -
WtP 89.2 98.2 96.5 - - - - 89.4 - 89.8 - - 94.1 98.2 -
WtPT 89.3 97.9 96.4 - - - - 92.0 - 87.3 - - 92.9 98.4 -
WtPPUNCT 97.3 99.6 99.0 - - - - 98.8 - 93.6 - - 97.0 99.4 -
SaT+LoRA 97.5 98.2 98.9 - - - - 96.5 - 90.9 - - 96.0 99.1 -
OPUS100 SpaCyM - 76.5 76.9 83.1 93.2 38.7 32.8 86.5 87.6 55.6 0.0 6.4 93.0 95.0 4.9
Ersatz - 77.0 77.6 - - - - - - - - - - - -
Llama 38B - 82.8 83.1 85.7 94.1 80.9 49.3 88.1 91.0 78.8 16.3 39.7 93.7 95.3 26.7
Command R - 70.3 70.4 64.0 48.2 0.7 26.6 65.2 67.9 65.1 17.1 23.3 74.8 71.5 14.4
SaT - 82.6 82.6 88.5 93.0 77.6 73.5 89.9 85.8 62.2 74.1 69.8 92.1 94.8 69.5
SaT+SM - 90.0 89.7 91.6 95.4 85.9 90.1 93.3 94.1 85.1 89.4 82.0 94.8 95.8 81.0
NLTK - - - - - 80.2 - - - - - - 93.4 94.5 -
PySBD - - - - - - - 86.2 - - 27.4 - 18.2 - -
SpaCyDP - 78.9 - - 82.2 - - - - - - - 92.4 - -
WtP - 76.5 78.0 89.1 92.3 80.0 80.7 88.5 87.0 60.6 68.9 68.9 91.5 94.2 55.6
WtPT - 84.1 85.6 91.5 92.4 81.7 - 88.6 87.9 80.4 74.3 68.7 - 94.3 62.2
WtPPUNCT - 90.1 91.5 95.3 95.6 86.5 - 93.5 94.0 88.4 82.2 74.3 - 96.1 77.3
SaT+LoRA - 92.6 92.8 95.4 95.5 86.5 - 94.9 93.8 89.2 86.7 77.3 - 96.3 79.6
Ersatz SpaCyM - 93.3 98.6 - - - - - - - - - - - -
Ersatz - 95.0 98.7 - - - - - - - - - - - -
Llama 38B - 94.9 98.8 - - - - - - - - - - - -
Command R - 78.3 82.7 - - - - - - - - - - - -
SaT - 96.3 97.5 - - - - - - - - - - - -
SaT+SM - 98.3 99.3 - - - - - - - - - - - -
NLTK - - - - - - - - - - - - - - -
PySBD - - - - - - - - - - - - - - -
SpaCyDP - 74.9 - - - - - - - - - - - - -
WtP - 96.5 96.9 - - - - - - - - - - - -
WtPT - 96.4 97.2 - - - - - - - - - - - -
WtPPUNCT - 99.2 99.3 - - - - - - - - - - - -
SaT+LoRA - 98.4 99.4 - - - - - - - - - - - -
Table 30: Sentence segmentation test F1 scores on languages la-pa.
Model pl ps pt ro ru si sk sl sq sr sv ta te tg th
UD SpaCyM
Ersatz 97.4 - - 98.3 77.7 - - - - - - 90.5 - - -
Llama 38B 99.3 - 93.5 93.5 88.5 - 90.6 98.6 100.0 96.5 95.5 96.7 - - 81.4
Command R 91.4 - 72.2 59.6 61.2 - 73.5 75.4 92.0 82.0 72.8 0.0 - - 12.4
SaT 95.3 - 96.6 73.2 82.7 - 94.8 96.8 100.0 98.0 95.0 99.5 - - 72.7
SaT+SM 99.2 - 97.2 99.0 92.4 - 96.5 99.2 99.1 99.4 96.2 99.5 - - 71.2
NLTK 97.2 - 91.9 - 78.3 - - - - - 93.9 - - - -
PySBD 84.8 - - - 67.9 - 86.2 - - - - - - - -
SpaCyDP 98.5 - 98.1 95.3 80.3 - - - - - 87.0 - - - -
WtP 94.4 - 95.9 80.9 84.9 - 96.0 95.7 100.0 97.9 94.7 96.9 - - 67.3
WtPT 95.5 - 95.6 93.5 86.8 - 95.8 96.2 - 98.2 95.0 97.7 - - -
WtPPUNCT 99.3 - 98.4 99.4 93.1 - 98.1 99.1 - 99.8 96.5 100.0 - - -
SaT+LoRA 98.9 - 97.8 99.5 92.7 - 96.9 99.2 - 99.7 96.5 99.5 - - -
OPUS100 SpaCyM 92.0 2.4 91.6 91.3 75.2 75.4 91.6 92.6 92.6 94.6 93.1 36.6 64.1 69.5 21.7
Ersatz 92.1 1.8 - 92.8 68.6 - - - - - - 45.3 - - -
Llama 38B 93.8 26.1 94.1 94.9 89.3 76.9 94.3 93.8 92.3 94.6 94.3 45.3 65.9 76.1 66.0
Command R 74.4 7.2 80.8 70.4 88.1 8.6 74.8 63.8 59.3 67.2 76.9 1.9 36.1 54.9 10.6
SaT 92.4 68.6 91.7 91.1 82.8 79.2 91.4 91.6 89.7 94.0 91.3 60.5 76.5 79.6 68.0
SaT+SM 95.6 85.6 93.8 95.9 84.1 85.4 95.9 95.3 95.7 95.9 95.0 75.0 86.8 92.3 72.9
NLTK 92.5 - 92.2 - 75.8 - - - - - 92.5 - - - -
PySBD 17.5 - - - 64.9 - 29.5 - - - - - - - -
SpaCyDP 92.9 - 90.8 91.9 75.4 - - - - - 90.2 - - - -
WtP 91.8 63.3 89.9 88.5 80.6 79.5 89.7 91.3 88.7 93.8 90.4 64.3 77.4 80.1 66.6
WtPT 92.2 70.4 91.4 89.0 - 80.1 92.4 92.7 89.9 94.3 92.5 65.7 77.5 82.7 69.7
WtPPUNCT 95.6 76.1 95.3 96.7 - 85.4 95.9 95.0 95.5 96.4 95.8 75.4 83.6 91.0 71.3
SaT+LoRA 96.0 77.6 95.4 97.3 - 85.9 96.5 95.3 95.5 96.1 95.9 80.8 85.6 92.1 73.7
Ersatz SpaCyM 93.3 94.3 - 94.8 93.2 - - - - - - 67.9 - - -
Ersatz 94.9 93.7 - 96.0 94.2 - - - - - - 95.2 - - -
Llama 38B 95.4 96.3 - 97.7 97.0 - - - - - - 97.0 - - -
Command R 70.1 70.7 - 71.3 80.2 - - - - - - 1.3 - - -
SaT 93.5 92.4 - 97.5 97.2 - - - - - - 97.6 - - -
SaT+SM 98.3 82.7 - 99.1 98.8 - - - - - - 98.5 - - -
NLTK 94.0 - - - 93.7 - - - - - - - - - -
PySBD 45.7 - - - 55.2 - - - - - - - - - -
SpaCyDP 94.5 - - 94.4 94.1 - - - - - - - - - -
WtP 94.6 83.7 - 97.5 97.5 - - - - - - 94.1 - - -
WtPT 92.8 91.0 - 96.9 97.6 - - - - - - 94.7 - - -
WtPPUNCT 97.7 95.9 - 99.4 99.4 - - - - - - 97.8 - - -
SaT+LoRA 98.1 96.1 - 99.3 98.8 - - - - - - 98.1 - - -
Table 31: Sentence segmentation test F1 scores on languages pl-th.
Model tr uk ur uz vi xh yi yo zh zu
UD SpaCyM 97.5 93.9 0.0 - 96.0 - - 79.2 0.0 -
Ersatz 96.8 - - - - - - - 89.3 -
Llama 38B 97.5 94.9 97.0 - 98.5 - - 89.8 95.8 -
Command R 61.8 62.1 82.3 - 92.4 - - 66.1 91.8 -
SaT 96.3 92.7 97.7 - 90.8 - - 77.0 94.5 -
SaT+SM 98.6 98.3 99.3 - 99.5 - - 89.6 98.6 -
NLTK 93.2 - - - - - - - - -
PySBD - - 99.2 - - - - - 98.9 -
SpaCyDP - 96.5 - - - - - - 99.0 -
WtP 95.9 92.0 91.7 - 88.5 - - 83.5 97.9 -
WtPT 95.6 92.1 95.8 - 93.7 - - - 98.0 -
WtPPUNCT 98.4 98.6 99.5 - 99.7 - - - 99.9 -
SaT+LoRA 98.5 98.1 99.5 - 99.4 - - - 99.3 -
OPUS100 SpaCyM 93.6 89.2 29.4 63.6 90.1 64.6 4.1 27.2 0.0 25.4
Ersatz 92.7 - - - - - - - 54.7 -
Llama 38B 93.8 91.1 47.6 67.4 92.9 71.7 13.6 37.5 55.7 42.5
Command R 72.4 79.8 26.6 44.8 77.2 56.1 8.3 26.6 58.7 34.9
SaT 93.2 88.3 51.3 76.3 91.1 80.3 61.1 67.3 55.6 82.1
SaT+SM 94.7 93.5 60.5 84.9 94.7 89.6 89.0 52.3 77.5 93.3
NLTK 93.4 - - - - - - - - -
PySBD - - 31.4 - - - - - 69.0 -
SpaCyDP - 89.8 - - - - - - 68.2 -
WtP 92.8 88.2 53.0 76.4 90.1 77.2 73.0 75.4 80.5 72.7
WtPT 93.1 89.0 50.7 78.9 90.3 80.7 73.9 - 76.6 83.1
WtPPUNCT 95.3 94.3 66.3 85.0 94.5 89.8 80.7 - 88.8 90.6
SaT+LoRA 95.4 94.7 72.0 87.6 94.8 90.8 86.6 - 90.5 92.0
Ersatz SpaCyM 95.2 - - - - - - - 0.0 -
Ersatz 96.2 - - - - - - - 87.4 -
Llama 38B 94.3 - - - - - - - 94.5 -
Command R 71.8 - - - - - - - 70.5 -
SaT 93.5 - - - - - - - 84.1 -
SaT+SM 97.5 - - - - - - - 90.5 -
NLTK 92.6 - - - - - - - - -
PySBD - - - - - - - - 92.7 -
SpaCyDP - - - - - - - - 95.9 -
WtP 92.8 - - - - - - - 93.7 -
WtPT 93.0 - - - - - - - 93.4 -
WtPPUNCT 98.3 - - - - - - - 97.9 -
SaT+LoRA 98.2 - - - - - - - 95.0 -
Table 32: Sentence segmentation test F1 scores on languages tr-zu.